### v1

# suppReferences

Abstract
Misinformation undermines public trust in science and democracy, particularly on social media where inaccuracies can spread rapidly. Experts and laypeople have shown to be effective in correcting misinformation by manually identifying and explaining inaccuracies. Nevertheless, this approach is difficult to scale, a concern as technologies like large language models (LLMs) make misinformation easier to produce. LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce plausible but false content and references, and limitations in addressing multimodal information. To address these issues, we propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information. By retrieving contextual evidence and refutations, Muse can provide accurate and trustworthy explanations and references. It also describes visuals and conducts multimodal searches for correcting multimodal misinformation. We recruit fact-checking and journalism experts to evaluate corrections to real social media posts across 13 dimensions, ranging from the factuality of explanation to the relevance of references. The results demonstrate Muse’s ability to correct misinformation promptly after appearing on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality corrections from laypeople by 29%. This work underscores the potential of LLMs to combat real-world misinformation effectively and efficiently.
Introduction
Misinformation—including misinformed, inaccurate, and misleading information—has far-reaching and detrimental effects on individuals and society[1, 2, 3, 4, 5]. It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[6, 3, 7]. Evidence has demonstrated that election misinformation helped fuel the January 6th attack on the U.S. Capitol, where five people died and more than 100 police officers were injured[8]. Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10]. The result is tragic; an analysis showed that COVID-19 vaccines could have prevented at least 232,000 of COVID-19-associated deaths between May 2021 and September 2022 in the U.S. alone[11]. Concerns over misinformation on social media have been particularly significant[12, 13, 14], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[15, 16].
Fortunately, high-quality and prompt misinformation correction, which broadly includes identifying what part(s) of content is (or is not) misinformation and explaining why that part of the content is misinformed, inaccurate, or misleading, has effectively reduced people’s false beliefs[17, 4, 18]. To this end, domain experts (e.g., on FactCheck.org) have been instrumental in providing accurate, detailed, and well-written explanations of whether and why a claim can be inaccurate with references. Subsequently, other research has shown that groups of laypeople can also identify and produce effective explanations to correct misinformation if their assessments are strategically aggregated[18, 19, 20]. Nevertheless, combating misinformation in the digital and generative artificial intelligence (AI) era cannot solely rely on human intelligence[21, 22]. The number of active fact-checking sites has remained stagnant since 2022[23]. Even though crowdsourcing is more scalable than expert-led activities, keeping pace with millions of posts daily on social media is impossible. As Brandolini’s law indicates[24], correcting misinformation is laborious—often requiring one to search and evaluate the credibility of related articles with evidence completing or refuting the claims in the given piece and write responses as justifications—whereas creating and spreading misinformation is easy. The absence of sufficient coverage markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[25].
While generative AI models such as LLMs (e.g., OpenAI’s GPT series[26]) raise concerns that they facilitate the creation of misinformation, they also potentially make scaling up the correction of misinformation possible. Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand. Indeed, LLMs have revolutionized the field of AI and presented remarkable capabilities across domains and tasks[27, 28, 29]. However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[30]. For instance, GPT-4[26] (as of March 2023) and MisinfoCorrect[31] lack access to timely updated knowledge and are thereby ill-equipped to combat misinformation on emerging topics. They either do not provide or “hallucinate” references, which can be fabricated or irrelevant[30, 32]. A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[33, 34, 35, 36, 37, 38, 39, 40]. However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating inaccuracies and backfiring (i.e., reinforcing rather than reducing false beliefs)[34, 41, 42]. These LLMs, as well as MisinfoCorrect[31], also struggle with counteracting multimodal misinformation due to their nonacceptance of visual inputs. Finally, understanding the quality of a correction necessitates a comprehensive human evaluation[30]. However, prior uses of LLMs for misinformation correction were evaluated by non-experts on limited aspects of their generated responses (e.g., the general quality of responses[31])[40], which is insufficient to comprehensively understand their performance in identifying and explaining inaccuracies, generated text, and references.
In this article, we propose Muse, a scalable approach for multimodal misinformation correction. Muse makes use of an LLM and augments it with the ability to handle visuals, access timely, factual, and credible knowledge on the web, retrieve evidence that refutes or contextualizes misinformation, and generate clear explanations with accurate and trustworthy references. We recruited experts in fact-checking and journalism to comprehensively evaluate corrections made by Muse, and compare them to baselines including GPT-4 and high-quality corrections written by the collective efforts of laypeople on social media. Our assessment measures the overall quality of corrections, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness of corrections, the relevance, factuality, fluency, coherence, and toxicity of generated text, and the reachability, relevance, and credibility of references. We find that overall Muse outperforms GPT-4 by 37% and high-quality corrections from laypeople by 29%, demonstrating Muse’s ability to correct both textual and visual misinformation across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine, promptly after appearing on social media.
Approach
Muse is designed to automatically generate a correction to a given piece of social media content that is potentially misinformation. As the input, the content can contain text with or without visual information. As the output, the correction should identify what part(s) of the content is (or is not) misinformation, explain why that part of the content is misinformation, and provide links as references. The general pipeline of Muse is illustrated in Fig. 1.
We start by introducing the details of Muse with a piece of text-only misinformation as the input. First, Muse generates queries based on an LLM from the misinformation (Fig. 1b; Methods). Each query acts as the input of a web search engine to access timely updated web content and obtain a list of web links directly relevant to the query (Fig. 1b; Methods). After scraping the content from these web links, Muse calculates their direct relevance to the misinformation and removes irrelevant web pages (Fig. 1b; Methods). Then, Muse determines the credibility of web pages by looking up their publishers’ factuality and bias ratings and selects pages with high factuality and minimal bias (Fig. 1c; Methods). Next, Muse leverages an LLM to extract sentences or paragraphs from each of the web pages as evidence, i.e., the text either explicitly refutes the misinformation or provides additional context (Fig. 1d; Methods). Finally, Muse generates a correction to the misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig. 1d; Methods).
Note that misinformation, especially when it comes to social media, often contains extraneous information that is irrelevant to a correction, including unverifiable opinions or emojis, such as the textual content of the misinformation in Fig. 1. Therefore, generating queries instead of simply using the text of the post improves web searches as a way of denoising misinformation; see example queries in Fig. 1b. Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs to be corrected, whereas generating one query may overlook some misinformed claims (Supplementary Fig. S1). Another concern may arise from filtering retrieved web pages based on how relevant the content is to the misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction. However, the increase in selected web pages drastically elevates the expense of Muse (Methods). We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating corrections (Supplementary Fig. S2). Moreover, as illustrated in Fig. 1c, Muse ranks the selected web pages by their publishers’ factuality and bias. It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation).
When the input is a piece of multimodal misinformation with textual and visual information, Muse first describes each image in text (Fig. 1a) so that the misinformation can be handled by any LLM on downstream tasks, including query generation (Fig. 1b), evidence extraction (Fig. 1d), and correction generation (Fig. 1d). Specifically, Muse augments image captioning models developed to describe an image in natural language with recognizing celebrities and optical characters based on an LLM (Fig. 1a; Methods). Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with additional features, including identification of celebrities and embedded text—information crucial for making accurate corrections[43]. For example, even a state-of-the-art image captioning model[44] may describe the visual misinformation in Fig. 1 as simply “list of banned books in Florida”. The description overlooks the listed titles of books that are essential for the visual misinformation’s correction (see more examples in Supplementary Fig. S3). In addition, Muse conducts multimodal search on the web and computes multimodal relevance to filter out irrelevant web pages (Fig. 1b; Methods).
Evaluation
Muse’s evaluation was based on X Community Notes data. Community Notes (communitynotes.twitter.com, formerly Twitter Birdwatch) was launched in 2021 to empower people on X to collaboratively add context to potentially misleading posts. It has been shown to effectively reduce the spread of misinformation[20]. Community Notes operates by enabling people on X, often laypeople, to identify whether a tweet is potentially misleading, write free-response corrections of any tweet, and assess whether other people’s corrections are helpful or unhelpful. The platform strategically aggregates people’s assessments with a transparent algorithm such that a correction has a higher helpfulness score if more people with diverse backgrounds (e.g., with different political ideologies) agree that the correction was helpful[20]. Besides canceling out inaccurate judgments made by some individuals, the algorithm has been shown to empirically address the “tyranny of the majority” and coordinated attempts at gaming[20, 19]. A correction with a sufficiently high helpfulness score is then displayed on the corresponding tweet on X and visible to the public[20] (Methods; Supplementary Fig. S4).
We selected all potentially misleading tweets from Community Notes that have at least two corrections, where one correction has a high helpfulness score and the other has an average helpfulness score (Methods). We further generated corrections to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods). Experts in fact-checking and journalism were recruited, onboarded, and asked to evaluate the quality of corrections by various approaches to the same tweet (Methods). Experts were blinded to which approach had generated each correction. The evaluation contains 13 specific criteria, covering how well the correction identifies and explains inaccuracies, the quality of generated text, and the quality of references (Methods). It also contains the overall quality of correction by taking all 13 evaluation criteria into account (Methods).
Our primary finding is that the overall quality of Muse-generated corrections is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Fig. 2a). The overall quality of Muse-generated corrections has an average score of 8.1 out of 10, 29% higher than laypeople’s high-helpfulness corrections (mean: 6.3; $p=3\times 10^{-48}$, by Mann-Whitney U test unless otherwise specified; $N=464$), 37% higher than GPT-4-generated corrections (mean: 5.9; $p=4\times 10^{-42}$; $N=464$), and 56% significantly higher than laypeople’s average-helpfulness corrections (mean: 5.2; $p=5\times 10^{-81}$; $N=462$). Despite statistical insignificance between the overall quality of laypeople’s high-helpfulness corrections and GPT-4-generated corrections ($p=0.4$; $N=464$), the overall quality of GPT-4-generated corrections has the highest variability, and GPT-4 generates more corrections with extremely low quality. The standard deviation of the overall quality of GPT-4-generated corrections is 2.7, vs only 2.0 for Muse and laypeople. 10% of GPT-4’s generated corrections have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness corrections, 3% for laypeople’s high-helpfulness corrections, and 2% for Muse-generated corrections. Note that laypeople’s corrections were created on average 14 hours after the tweet was posted on social media. Here, Muse only retrieved web pages published before the tweet was posted (Methods).
Examining specific components of correction quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in identifying and explaining inaccuracies (Fig. 2b-f). Experts assessed that Muse-generated corrections more explicitly identify and explain where and why the tweets are inaccurate than GPT-4’s and laypeople’s high-helpfulness corrections (Fig. 2b). 89% of Muse’s generated corrections explicitly identify and explain inaccuracies, 16% more than GPT-4-generated corrections, 29% more than laypeople’s high-helpfulness corrections, and 43% more than laypeople’s average-helpfulness corrections (Fig. 2b). As for identifying where the tweets are inaccurate, we found that Muse more comprehensively identifies the tweets’ inaccuracies with fewer mistakes (i.e., claiming where the tweets should be inaccurate as accurate or vice versa) than GPT-4 and laypeople who produce even high-helpfulness corrections (Fig. 2c-d). 91% of Muse’s generated corrections have at least one accurate identification without any mistake, 11% more than laypeople’s high-helpfulness corrections, 19% more than GPT-4-generated corrections, and 26% more than laypeople’s average-helpfulness corrections (Fig. 2c). Muse has 61% of generated corrections accurately identifying all the tweets’ inaccuracies, vs GPT-4 has 38%, laypeople who produce high-helpfulness corrections have 26%, and laypeople who produce average-helpfulness correction have 17% only (Fig. 2d). Furthermore, Muse explains why the tweets are inaccurate more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness corrections (Fig. 2e-f). 70% of corrections by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness corrections, 47% by GPT-4, and 37% for laypeople’s average-helpfulness corrections only (Fig. 2e). Meanwhile, the average informativeness score of Muse-generated corrections is 8.4, 18% higher than GPT-4-generated corrections, 29% higher than laypeople’s high-helpfulness corrections, and 50% higher than laypeople’s average-helpfulness corrections (Fig. 2f).
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in the quality of generated text (Fig. 2g-k). Muse, which can augment GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals, exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text in correcting misinformation compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=464$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness corrections by laypeople and additionally less toxic than the text of average-helpfulness corrections by laypeople ($p=4\times 10^{-12}$; $N=462$). In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness corrections by laypeople, and 43% higher than the text of average-helpfulness corrections by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness corrections by laypeople and 45% for GPT-4-generated text (Fig. 2h). Muse has 99% generated text without any mistake in the use of English, vs laypeople who produce even high-helpfulness corrections have 89% such text (Fig. 2i). Around 91% of Muse-generated text is highly coherent and logical, vs 76% and 61% for the text of high- and average-helpfulness corrections by laypeople (Fig. 2j). Around 99% of Muse-generated text is not biased, impolite, and provoking, 12% higher than that of average-helpfulness corrections by laypeople (Fig. 2k).
Additionally, results reveal that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in the quality of references (Fig. 2l-n). First, GPT-4 hallucinates references frequently. 49% of its links result in “page-not-found” errors (Fig. 2l), and only 76% of reachable links support or are relevant to the generated text (Fig. 2m). Muse significantly reduces such hallucinations with nearly 100% links being reachable (Fig. 2l) and 96% reachable links being relevant to the generated text (Fig. 2m). Meanwhile, Muse’s references are more credible than the references offered in even high-helpfulness corrections by laypeople ($p=4\times 10^{-11}$; $N=744$; Fig. 2n).
Furthermore, we observed that the quality of Muse-generated corrections to text-only and multimodal (textual and visual) misinformation is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Fig. 3a; Supplementary Fig. S5-S6). The quality of Muse-generated corrections to text-only tweets is 29% higher than GPT-4-generated corrections ($p=6\times 10^{-28}$; $N=310$), 33% higher than laypeople’s high-helpfulness corrections ($p=10^{-44}$; $N=310$), and 65% higher than laypeople’s average-helpfulness corrections ($p=10^{-67}$; $N=310$). The quality of Muse-generated corrections to multimodal tweets is generally 21% higher than laypeople’s high-helpfulness corrections ($p=3\times 10^{-8}$; $N=154$), 39% higher than laypeople’s average-helpfulness corrections ($p=10^{-16}$; $N=154$), and 56% higher than GPT-4 ($p=6\times 10^{-18}$; $N=152$).
Finally, we found that the quality of Muse-generated corrections to misinformation across domains including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Methods; Fig. 3b; Supplementary Fig. S7-S11). The quality of Muse-generated corrections is 35% (political and international affair; $p=10^{-16}$; $N=160$), 47% (economy and business; $p=4\times 10^{-11}$; $N=76$), 44% (crime and law; $p=3\times 10^{-8}$; $N=76$), 31% (social issues and human rights; $p=0.0002$; $N=60$), and 27% (health and medicine; $p=3\times 10^{-5}$; $N=48$) higher than GPT-4-generated corrections. The quality of Muse-generated corrections is 33% (political and international affair; $p=10^{-21}$; $N=160$), 27% (economy and business; $p=10^{-8}$; $N=76$), 28% (crime and law; $p=3\times 10^{-9}$; $N=76$), 38% (social issues and human rights; $p=7\times 10^{-10}$; $N=60$), and 25% (health and medicine; $p=4\times 10^{-5}$; $N=48$) higher than even high-helpfulness corrections by laypeople.
Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[21, 22], our work demonstrates the potential of LLMs to improve the online information ecosystem by correcting misinformed, inaccurate, and misleading information[2, 3]. Correcting misinformation, particularly on social media, is complex. It requires comprehending multimodal content, understanding context involving emerging topics or events, accurately and comprehensively identifying stated inaccuracies that are subtly different from the facts, and writing easy-to-understand, precise, and informative explanations supported by trustworthy references. We propose Muse and demonstrate the high quality of its automatically generated corrections to misinformation on social media. Muse augments existing powerful LLMs (e.g., GPT-4[26]) with the capabilities of addressing visuals, accessing up-to-date knowledge, and finding accurate references nonparametrically (i.e., without the need to train or fine-tune a machine learning model) such that Muse can be easily and cheaply updated, especially compared to parametric models[35]. Results further validate that addressing visual information is crucial in correcting multimodal misinformation, where GPT-4 struggles but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining inaccuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4.
We provide Muse as a solution to assist social media users and platforms in accurately, scalably, promptly, and transparently correcting misinformation. Muse is end-to-end and thereby simple to use. Meanwhile, our results reveal that Muse-generated corrections have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness corrections in correcting misinformation across modalities and domains (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated corrections show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating corrections with the least toxic text and most credible references[31, 41, 42] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[25] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. These high-quality corrections by Muse can generally be obtained two minutes after misinformation appears on social media (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify corrections themselves (Fig 2,3f).
Muse’s corrections cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the complexity of this task and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[18] to identify whether a new article’s headline and lede contains misinformation without explaining the inaccuracies.
This work also faces the following limitations. First, although Muse is capable of addressing and correcting multimodal misinformation with text and images, it cannot accept video inputs. Second, we focus on correcting misinformation in a single language, English, one of the most spoken languages in the world, whereas misinformation has become a global issue and exists in various high- and low-resource languages [45]. Third, we evaluated Muse using real-world misinformation on a single social media platform, X, as its Community Notes system provides publicly accessible large-scale data and a transparent state-of-the-art aggregation algorithm, enabling us to compare Muse’s corrections with the high- and average-helpfulness corrections made by social media users. X is also a popular social media platform, and one where more than half of users consume news regularly[46] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one other LLM, GPT-4[26], which can also be seen as an ablation study, since Muse augments GPT-4 (Methods). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[47, 48].
Conclusion
We proposed Muse, a nonparametric LLM for scaling up high-quality and prompt corrections of misinformation on social media, that can integrate textual and visual information, access timely updated knowledge, and generate easy-to-understand corrections with accurate and trustworthy references. Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who produce high-helpfulness corrections on X Community Notes in identifying and explaining inaccuracies, generating high-quality text, providing high-quality references, and producing the highest quality of corrections overall. Muse also excels in correcting text and multimodal (textual and visual) misinformation and misinformation across politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine. This study demonstrates the potential of LLMs in correcting misinformation online effectively, scalably, promptly, and transparently.
Acknowledgements
We would like to thank Hacks/Hackers for advertising the study and helping with the recruitment. We also thank members of the UW Social Futures Lab and Behavioral Data Science Group for their suggestions and feedback. This work was supported in part by the National Science Foundation’s Convergence Accelerator program under Award No. 49100421C0037.
Author contributions
X.Z., A.X.Z., and T.A. designed the study. X.Z. led, and A.S. assisted in developing the model. A.S. built the web interface for evaluation. X.Z. prepared and analyzed the data. All authors contributed to writing the manuscript. A.X.Z. and T.A. supervised the study.
Competing interests
The authors declare no competing interests.
Additional information
Supplementary Information is available for this paper.
Correspondence and requests for materials should be addressed to Tim Althoff (althoff@cs.washington.edu) and Amy X. Zhang (axz@cs.uw.edu).
Supplementary Materials
List of supplementary materials
Table S1
Figures S1 to S27
Table S1: Definitions and statistics of factuality and bias categories offered by Media Bias/Fact Check (mediabiasfactcheck.com).
Category
Definition
# of sources
Factuality:
- Very high:
The source is consistently factual, relies on credible information, promptly corrects errors, and has never failed any fact checks in news reporting or opinion pieces.
118
- High:
The source is mostly factual and uses mostly credible, low-biased, or high-factual sources. It corrects errors quickly and has failed only one news fact check and up to two op-ed fact checks.
2,313
- Mostly factual:
The source is generally accurate but may have a few uncorrected fact-check failures. It can fail up to three op-ed fact checks, especially if it is a low-volume site. While it may use biased sources occasionally, it mostly links to factual content. It is usually pro-science but may sometimes use misleading wording or offer alternative viewpoints. It is reasonably transparent and trustworthy most of the time, but caution is advised.
326
- Mixed:
The source may rely on improper sourcing or link to other biased or mixed-factual sources. It often has multiple failed fact checks and does not correct false information or lacks transparency, including the absence of a disclosed mission statement or ownership details. Sources rejecting established scientific consensus on issues such as climate change or vaccines will receive this rating or lower. Sources identified as overt propaganda or designated as hate groups by reputable third-party evaluators will receive this rating or lower due to their inherent bias and potential spread of misleading information.
1,437
- Low:
The source is often unreliable and should be fact-checked for fake news, conspiracy theories, and propaganda.
677
- Very low:
The source is almost always unreliable and should always be fact-checked for intentional misinformation.
252
Total: 5,123
Bias:
- Least biased:
The source has minimal bias and uses very few loaded words (i.e., wording that attempts to influence an audience by using an appeal to emotion or stereotypes). It is factual and usually sourced.
1,054
- Left-center:
The source has a slight to moderate liberal bias. It often publishes factual information that utilizes loaded words to favor liberal causes. It is generally trustworthy for information but may require further investigation.
850
- Right-center:
Similar to the definition of left-center bias but replacing liberal with conservative.
492
- (Extremely) left:
The source is moderately to strongly biased toward liberal causes through story selection or political affiliation. It may utilize strong loaded words, publish misleading reports, and omit reporting of information that may damage liberal causes. It may be untrustworthy.
402
- (Extremely) right:
Similar to the definition of (extremely) left bias but replacing liberal with conservative.
314
- Pro-science:
The source consists of legitimate science or is evidence-based through the use of credible scientific sourcing. Legitimate science follows the scientific method, is unbiased, and does not use emotional words. The source also respects the consensus of experts in the given scientific field and strives to publish peer-reviewed science. It may have a slight political bias but adheres to scientific principles.
189
- Conspiracy- pseudoscience:
The source may publish unverifiable information not always supported by evidence. It may be untrustworthy for credible or verifiable information, so fact-checking and further investigation are recommended on a per-article basis when obtaining information from it.
433
- Questionable:
The source exhibits one or more of the following: extreme bias, consistent promotion of propaganda or conspiracies, poor or no sourcing of credible information, a complete lack of transparency, or fake news (i.e., the deliberate attempt to publish hoaxes or disinformation for profit or influence). It may be very untrustworthy and should be fact-checked on a per-article basis.
1,390
- Satire:
The source exclusively uses humor, irony, exaggeration, or ridicule to expose and criticize people’s stupidity or vices, particularly in the context of contemporary politics and other topical issues. It does not attempt to deceive.
148
Total: 5,272
Figure S1: Examples that show that generating multiple queries helps decompose a post, which may have multiple claims that each need to be corrected, whereas generating one query may overlook some misinformed, inaccurate, or misleading claims. Bold text: the misinformed, inaccurate, or misleading claims overlooked when generating one query but captured when generating more than one query.
Figure S2: Examples that show how retrieved web pages with relatively low relevance to misinformation can promote LLM (in this case, GPT-4) hallucinations when generating corrections.
Figure S3: Examples of informative image captions, which augment image captions with names of visually represented celebrities and embedded text (see Methods for implementation details).
Figure S4: An example of a high-helpfulness correction from Community Notes displayed on the corresponding tweet on X and visible to the public.
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S5: Expert evaluation results for text misinformation (#=155).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S6: Expert evaluation results for multimodal (textual and visual) misinformation (#=77).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S7: Expert evaluation results for political and international-affair misinformation (#=80).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S8: Expert evaluation results for economic and business misinformation (#=38).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S9: Expert evaluation results for crime and law misinformation (#=38).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S10: Expert evaluation results for social-issue and human-right misinformation (#=30).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S11: Expert evaluation results for health and medicine misinformation (#=24).
Figure S12: Images as examples used for informative image captioning with in-context learning.
1Describe an image in an informative way. Your description should be only based on the given {short caption}, {name of each person}, and {raw text}. If the image is from social media, you should start with "A screenshot of". If the image is a quote from someone, you should start with "A quote from" followed by this person’s name if there is any, then by the quoted text. If the image is an article, you should start with "An article". If the image is a photo, you should start with "A photo of". If the image is a map, you should start with "A map of". {raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.2short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}3name of each person: {Fran Lebowitz}4raw text: {"In real life, I assure you, there is no such thing as algebra."}5image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}6short caption: {two men in suits}7name of each person: {Jim Caviezel, Michael Emerson}8raw text: {}9image description: {A photo of Jim Caviezel and Michael Emerson in suits}10short caption: {robbie leems on twitter}11name of each person: {}12raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother! 1 2 3,790}13image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}14short caption: {a moose}15name of each person: {}16raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession. finance.yahoo.com Could the stock market power through a recession? ’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}17image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession? ’This would be rare.’" with a picture of a moose. The post was posted at 09:57 22/4/2023.}18short caption: {a person pouring tea into a cup}19name of each person: {}20raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}21image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup. It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}22short caption: {two people standing next to each other with the words love is blind}23name of each person: {Nick Lachey}24raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}25image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other. It was written by Hayley Miller and published by MSNBC.}26short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}27name of each person: {Rahul Kanwal}28raw text: {How engaged are the most-followed journalists on Twitter? Percentage of tweets from each journalist that are at-replies BDUTT 64% RealMikeWilbon 38% 34% ErinAndrews 31% stephenasmith 24% andersoncooper rahulkanwal 13% 13% ninagarcia 5% sardesairajdeep 1% BillSimmons maddow 1% 0% 10% 20% 30% 40% 50% 60% chart by mattmaldre.com}29image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies. The chart was made by mattmaldre.com.}30short caption: {a graph showing the global defense budget by region}31name of each person: {}32raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}33image description: {A graph showing the global defense budget by region. It is from Aviation Week.}34short caption: {[IMAGE_CAPTION]}35name of each person: {[CELEBRITIES]}36raw text: {[OCR]}37image description:
Figure S13: LLM prompt for informative image captioning.
Figure S14: Distribution of helpfulness scores of laypeople’s corrections in X Community Notes. All: All laypeople’s corrections in Community Notes. Average helpfulness: Laypeople’s corrections in Community Notes identified with average helpfulness and used in our study. High helpfulness: Laypeople’s corrections in Community Notes identified with high helpfulness and used in our study. For $x\pm y$, $x$: mean, $y$: standard deviation. Community Notes data are regularly updated; ours are up until February 12, 2023.
Figure S15: Distribution of creation times of laypeople’s corrections in X Community Notes used in our study. Median of the creation time of laypeople’s average-helpfulness corrections: 16 hours after the misinformation was posted. Median of the creation time of laypeople’s high-helpfulness corrections: 13 hours after the misinformation was posted.
Figure S16: Annotation instructions (page 1/7, continued on the next page).
Figure S17: Annotation instructions (page 2/7, continued on the next page).
Figure S18: Annotation instructions (page 3/7, continued on the next page).
Figure S19: Annotation instructions (page 4/7, continued on the next page).
Figure S20: Annotation instructions (page 5/7, continued on the next page).
Figure S21: Annotation instructions (page 6/7, continued on the next page).
Figure S22: Annotation instructions (page 7/7).
Figure S23: Annotation task page.
Figure S24: Impact of starting times of correcting misinformation on Muse’s performance. The simulated starting time for Muse (avg): Thirty minutes before the corresponding laypeople’s average-helpfulness correction was created (median: 16 hours after the corresponding misinformation was posted; Supplementary Fig. S15). The simulated starting time for Muse (high): Thirty minutes before the corresponding laypeople’s high-helpfulness correction was created (median: 13 hours after the corresponding misinformation was posted; Supplementary Fig. S15). The simulated starting time for Muse: The post time of the corresponding misinformation (i.e., 0 hours after the corresponding misinformation was posted).
Figure S25: Distribution of post times of misinformation in X Community Notes included in our analyses.
Figure S26: Impact of post times of misinformation on Muse’s performance.
(a) Overall quality of corrections.
(b) Quality of corrections in identifying and explaining inaccuracies.
(c) Quality of corrections in generated text.
(d) Quality of corrections in references.
[…]