### v3

# suppReferences

Abstract
Real-world misinformation can be partially correct and even factual but misleading. It undermines public trust in science and democracy, particularly on social media, where it can spread rapidly. High-quality and timely correction of misinformation that identifies and explains its (in)accuracies has been shown to effectively reduce false beliefs. Despite the wide acceptance of manual correction, it is difficult to be timely and scalable, a concern as technologies like large language models (LLMs) make misinformation easier to produce. LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce false content, and limitations in addressing multimodal information. We propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information. By retrieving evidence as refutation or supporting context, Muse identifies and explains (in)accuracies in a piece of content—not presupposed to be misinformation—with references. It also describes images and conducts multimodal searches to verify and correct multimodal content. Fact-checking experts evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts that may or may not be misleading. We propose and evaluate 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references. The results demonstrate Muse’s ability to promptly write high-quality responses to potential misinformation on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%. This work reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
Introduction
Misinformation, which broadly includes partially incorrect and factual but misleading content[1, 2], has far-reaching and detrimental effects on individuals and society[3, 1, 2, 4, 5]. It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[6, 2, 7]. Evidence has demonstrated that election misinformation helped fuel the January 6th attack on the U.S. Capitol, where five people died and more than 100 police officers were injured[8]. Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10]. The result is tragic, where COVID-19 vaccines could have prevented at least 232,000 of COVID-19-associated deaths between May 2021 and September 2022 in the U.S. alone[11]. Concerns over misinformation on social media have been particularly significant[12, 13, 14], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[15, 16].
Fortunately, high-quality and timely correction of misinformation, which identifies what part(s) of its content is or is not accurate and explains why that part of the content is (in)accurate with references, has been shown to effectively reduce the spread of misinformation and false beliefs[17, 18, 4, 19]. While domain experts (e.g., on FactCheck.org) and groups of laypeople (e.g., on X Community Notes, formerly Twitter Birdwatch) have played pivotal roles in correcting misinformation[19, 20, 17], keeping pace with massive social media posts is impossible. As Brandolini’s law indicates[21], correcting misinformation is laborious—often requiring one to search for related and trusted articles and write justifications—whereas creating and spreading misinformation is easy. As a consequence, a 88% of suspicious content on X did not receive any response, and 93% did not receive a high-quality response within the first hour (according to X/Twitter Community Notes as of February 2023; Supplementary Fig. S1). Even high-quality responses suffer from limited effectiveness, when they are created after rather than before initial bursts of attention to misinformation content[4, 14]. The absence of sufficient coverage also markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[22].
While generative AI models such as LLMs (e.g., OpenAI’s GPT series) raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible. Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand. Indeed, LLMs have revolutionized the field of AI and presented remarkable capabilities across domains and tasks[23, 24, 25]. However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[26]. For instance, GPT-4[27] (as of March 2023) and MisinfoCorrect[28] lack access to timely updated knowledge and are thereby ill-equipped to combat misinformation on emerging topics. They either do not provide or “hallucinate” references, which can be fabricated or irrelevant[26, 29]. A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[30, 31, 32, 33, 34, 35, 36, 37]. However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating falsehood and backfiring (i.e., reinforcing rather than reducing false beliefs)[31, 38, 39]. These LLMs, as well as MisinfoCorrect[28], also struggle with counteracting multimodal misinformation due to their nonacceptance of visual inputs. Finally, understanding the quality of a correction necessitates a comprehensive evaluation of fact-checking experts due to the complexity of real-world misinformation[26]. However, prior evaluation of LLMs was conducted by non-experts on limited aspects of their generated responses (e.g., the general quality of responses[28])[37], which is insufficient to comprehensively understand their performance in identifying and explaining (in)accuracies, generated text, and references.
In this article, we propose Muse, a scalable approach for multimodal misinformation correction. Muse makes use of an LLM and augments it with the ability to handle images, access timely and credible knowledge on the web, retrieve evidence that refutes or contextualizes the given content that may or may not be misinformation, and generate clear explanations with accurate and trustworthy references. Fact-checking experts comprehensively evaluate Muse-generated responses to real social media posts that potentially are misinformation, and compare them to baselines including GPT-4 and high-quality responses based on the collective efforts of laypeople. Our assessment measures the overall quality of a response, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness when it identifies and explains (in)accuracies, the relevance, factuality, fluency, coherence, and toxicity of generated text, and the reachability, relevance, and credibility of references. We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in effectively and promptly responding to potential misinformation. Results demonstrate Muse’s advance when the content is textual with or without images across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine.
Approach
Muse is designed to automatically respond to content that potentially is misinformation. In other words, the content might be inaccurate, partially accurate, or even factually accurate but misleading, all of which are misinformation, or fully accurate as non-misinformation. The content can contain text with or without visual information. The response should identify what part(s) of the content is (in)accurate, explain why that part of the content is (in)accurate, and provide links as references. The pipeline of Muse is illustrated in Fig. 1.
We start by introducing the details of Muse with a piece of text-only potential misinformation as the input. First, Muse generates queries based on an LLM from the potential misinformation (Fig. 1b; Methods). Each query acts as the input of a web search engine to access timely updated web content and obtain a list of web links directly relevant to the query (Fig. 1b; Methods). After scraping the content from these web links, Muse calculates their direct relevance to the potential misinformation and removes irrelevant web pages (Fig. 1b; Methods). Then, Muse determines the credibility of web pages by looking up their publishers’ factuality and bias ratings and selects pages with high factuality and minimal bias (Fig. 1c; Methods). Next, Muse leverages an LLM to extract text from each of the web pages as evidence. Such evidence can refute the potential misinformation, typically happening when it is misinformation with false claims, or provide additional context, which can demonstrate that the potential misinformation is accurate or part(s) of its claims are accurate (Fig. 1d; Methods). Finally, Muse generates a response to the potential misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig. 1d; Methods).
Note that content, especially content posted on social media, often contains extraneous information that does not need verification and is irrelevant to correction, including unverifiable opinions or emojis, such as the textual content of the false post in Fig. 1. Therefore, generating queries instead of simply using the post content improves web searches as a way of denoising the post content; see example queries in Fig. 1b. Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs verification or correction, whereas generating one query may overlook some of the claims and hence lead to not comprehensive identifications and explanations of (in)accuracies (Supplementary Fig. S2). Another concern may arise from filtering retrieved web pages based on how relevant the page content is to the potential misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction. However, the increase in selected web pages drastically elevates the expense of Muse (Methods). We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating responses (Supplementary Fig. S3). Moreover, as illustrated in Fig. 1c, Muse filters and ranks the selected web pages by their publishers’ factuality and bias. It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation) or gone through all the credible pages.
When the input is a piece of multimodal (textual and visual) content, Muse first generates textual description of each image (Fig. 1a) so that the content can be handled by any LLM on downstream tasks, including query generation (Fig. 1b), evidence extraction (Fig. 1d), and response generation (Fig. 1d). Specifically, Muse augments image captioning models developed to describe an image in natural language with recognizing celebrities and optical characters based on an LLM (Fig. 1a; Methods). Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with additional features, including identification of celebrities and embedded text—information crucial for making accurate verification and corrections[40]. For example, even a state-of-the-art image captioning model[41] may describe the visual misinformation in Fig. 1 as simply “list of banned books in Florida.” The description overlooks the listed titles of books that are essential for the visual content’s verification and correction (see more examples in Supplementary Fig. S4). In addition, Muse conducts multimodal search on the web and computes multimodal relevance to filter out irrelevant web pages (Fig. 1b; Methods).
Evaluation
Muse’s evaluation was based on X Community Notes data. Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has been shown to reduce the spread of misinformation[17]. Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies. A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supplementary Fig. S6). We included the tweets from Community Notes with at least two responses (#=247); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods). Though we do not presuppose the accuracy of the tweets in Muse’s design and evaluation, we found that more than half of the tweets are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods). We further generated responses to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods). Experts in fact-checking and journalism evaluated the quality of responses by various approaches to the same tweet (Methods); they were blinded to which approach had generated each response. The evaluation contains 13 specific criteria, covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references (Methods). It also contains the overall quality of a response by taking all 13 evaluation criteria into account (Methods).
Our primary finding is that the overall quality of Muse-generated responses is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig. 2a). The overall quality of Muse-generated responses has an average score of 8.1 out of 10, 29% higher than laypeople’s high-helpfulness responses (mean: 6.3; $p=3\times 10^{-48}$, by Mann-Whitney U test unless otherwise specified; $N=464$), 37% higher than GPT-4-generated responses (mean: 5.9; $p=4\times 10^{-42}$; $N=464$), and 56% significantly higher than laypeople’s average-helpfulness responses (mean: 5.2; $p=5\times 10^{-81}$; $N=462$). Despite statistical insignificance between the overall quality of laypeople’s high-helpfulness responses and GPT-4-generated responses ($p=0.4$; $N=464$), the overall quality of GPT-4-generated responses has the highest variability, and GPT-4 generates more responses with extremely low quality. The standard deviation of the overall quality of GPT-4-generated responses is 2.7, vs only 2.0 for Muse and laypeople. 10% of GPT-4’s generated responses have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness responses, 3% for laypeople’s high-helpfulness responses, and 2% for Muse-generated responses. Note that laypeople’s responses were created on average 14 hours after the tweet was posted on social media. Here, Muse only retrieved web pages published before the tweet was posted (Methods).
Examining specific components of response quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in identifying and explaining (in)accuracies (Fig. 2b-f). Experts assessed that Muse-generated responses more explicitly identify and explain where and why a tweet is (in)accurate than GPT-4’s and laypeople’s high-helpfulness responses (Fig. 2b). 89% of Muse’s generated responses explicitly identify and explain (in)accuracies, 16% more than GPT-4-generated responses, 29% more than laypeople’s high-helpfulness responses, and 43% more than laypeople’s average-helpfulness responses (Fig. 2b). As for identifying where a tweet is (in)accurate, we found that Muse more comprehensively identifies a tweet’s (in)accuracies with fewer mistakes—here, mistakes indicate falsely claiming where a tweet should be inaccurate as accurate or where a tweet should be accurate as inaccurate—than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2c-d). 91% of Muse’s generated responses have at least one correct identification without any mistake, 11% more than laypeople’s high-helpfulness responses, 19% more than GPT-4-generated responses, and 26% more than laypeople’s average-helpfulness responses (Fig. 2c). Muse has 61% of generated responses accurately identifying all the (in)accuracies in a tweet, vs GPT-4 has 38%, laypeople who produce high-helpfulness responses have 26%, and laypeople who produce average-helpfulness responses have 17% only (Fig. 2d). Furthermore, Muse explains (in)accuracies more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2e-f). 70% of responses by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness responses, 47% by GPT-4, and 37% for laypeople’s average-helpfulness responses only (Fig. 2e). Meanwhile, the average informativeness score of Muse-generated responses is 7.9, 32% higher than laypeople’s high-helpfulness responses, 36% higher than GPT-4-generated responses, and 65% higher than laypeople’s average-helpfulness responses (Fig. 2f).
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of generated text (Fig. 2g-k). Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=463$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$). In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness responses by laypeople and 45% for GPT-4-generated text (Fig. 2h). Almost all of Muse-generated text does not have any mistake in the use of English (Fig. 2i) and is not biased, impolite, and provoking (Fig. 2k), and 91% is highly coherent and logical (vs 76% and 61% for laypeople, Fig. 2j).
Additionally, results reveal that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of references (Fig. 2l-n). First, GPT-4 hallucinates references frequently. 49% of its links result in “page-not-found” errors (Fig. 2l), and only 76% of reachable links are relevant to the generated text (Fig. 2m). Muse significantly reduces such hallucinations with nearly 100% links being reachable (Fig. 2l) and 96% reachable links being relevant to the generated text (Fig. 2m). Meanwhile, Muse’s references are more credible than the references offered in even high-helpfulness responses by laypeople ($p=4\times 10^{-11}$; $N=744$; Fig. 2n).
Furthermore, we observed that the quality of Muse-generated responses to textual and multimodal (textual and visual) content that potentially is misinformation is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig. 3a; Supplementary Fig. S7-S8). The quality of Muse-generated responses to text-based tweets is 29% higher than GPT-4-generated responses ($p=6\times 10^{-28}$; $N=310$), 33% higher than laypeople’s high-helpfulness responses ($p=10^{-44}$; $N=310$), and 65% higher than laypeople’s average-helpfulness responses ($p=10^{-67}$; $N=310$). The quality of Muse-generated responses to multimodal tweets is generally 21% higher than laypeople’s high-helpfulness responses ($p=3\times 10^{-8}$; $N=154$), 39% higher than laypeople’s average-helpfulness responses ($p=10^{-16}$; $N=154$), and 56% higher than GPT-4 ($p=6\times 10^{-18}$; $N=152$).
Finally, we found that the quality of Muse-generated responses to potential misinformation across domains including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Methods; Fig. 3b; Supplementary Fig. S9-S13). The quality of Muse-generated responses is 35% (political and international affair; $p=10^{-16}$; $N=160$), 47% (economy and business; $p=4\times 10^{-11}$; $N=76$), 44% (crime and law; $p=3\times 10^{-8}$; $N=76$), 31% (social issues and human rights; $p=0.0002$; $N=60$), and 27% (health and medicine; $p=3\times 10^{-5}$; $N=48$) higher than GPT-4-generated responses. The quality of Muse-generated responses is 33% (political and international affair; $p=10^{-21}$; $N=160$), 27% (economy and business; $p=10^{-8}$; $N=76$), 28% (crime and law; $p=3\times 10^{-9}$; $N=76$), 38% (social issues and human rights; $p=7\times 10^{-10}$; $N=60$), and 25% (health and medicine; $p=4\times 10^{-5}$; $N=48$) higher than even high-helpfulness responses by laypeople.
Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[42, 43], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[1, 2]. Real-world misinformation can combine (in)accurate, factually accurate but misleading, and unverifiable claims. Identifying a piece of content as misinformation without further interventions, done by many previous AI-driven approaches[44], has limited impact on reducing misinformation’s detrimental effects. Showing the identification results without explanations might even increase false beliefs[45]. Correcting misinformation has been shown to reduce its spread and false beliefs, whose effectiveness is affected by quality, timeliness, and scalability[4, 22]. However, research on scaling up and accelerating misinformation correction—which may include automation—is still at the early stage. Existing AI models struggle to identify and correct misinformation, especially on social media, where misinformation is not restricted to narrow domains and spreads rapidly. Correcting misinformation requires comprehending content that can be multimodal and the context beyond it that often involves emerging events. It requires identifying what part(s) of the content is (in)accurate and explaining why that part of the content is (in)accurate with trustworthy references. We propose Muse and demonstrate the high quality of its automatically generated responses to social media posts that potentially are misinformation. Muse augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references. Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining (in)accuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4.
We provide Muse as a solution to assist social media users and platforms in accurately, scalably, promptly, and transparently responding to suspicious content. Muse is end-to-end and thereby simple to use. It is nonparametric, i.e., does not need training or fine-tuning an AI model, and thereby easily and cheaply updated, especially compared to parametric models[32]. Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities and domains (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated responses show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating responses with the least toxic text and most credible references[28, 38, 39] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[22] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. These high-quality responses by Muse can generally be obtained two minutes after suspicious content (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f).
Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the task’s complexity and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[19] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation.
This work also faces the following limitations. First, although Muse is capable of responding to multimodal misinformation with text and images, it cannot accept video inputs. Second, we only focus on English, one of the most spoken languages in the world. Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has been shown to reduce the spread of misinformation[17] and transparent. X is also a popular social media platform, and one where more than half of users consume news regularly[46] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one other LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (Methods). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[47, 48].
Conclusion
We proposed Muse, a nonparametric LLM to scale up and accelerate misinformation corrections on social media. Muse can integrate textual and visual information, access timely knowledge, and generate responses that use natural languages and have accurate and trustworthy references. Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who produce high-helpfulness responses on X Community Notes in identifying and explaining inaccuracies, generating high-quality text, providing high-quality references, and producing the highest quality of responses overall. Muse also excels when the responded content is textual or multimodal (textual and visual) and when it is related to politics and international affairs, economy and business, crime and law, social issues and human rights, or health and medicine. This study demonstrates the potential of LLMs in responding to online misinformation effectively, scalably, promptly, and transparently.
Acknowledgements
We would like to thank Hacks/Hackers for advertising the study and helping with the recruitment. We also thank members of the UW Social Futures Lab and Behavioral Data Science Group for their suggestions and feedback. This work was supported in part by the National Science Foundation’s Convergence Accelerator program under Award No. 49100421C0037. T.A. and A.S. were supported in part by the Office of Naval Research (#N00014-21-1-2154), NSF IIS-1901386, and NSF CAREER IIS-2142794.
Author contributions
X.Z., A.X.Z., and T.A. designed the study. X.Z. led, and A.S. assisted in developing the model. A.S. built the web interface for evaluation. X.Z. prepared and analyzed the data. X.Z. drafted and all authors revised the manuscript. A.X.Z. and T.A. supervised the study.
Competing interests
The authors declare no competing interests.
Additional information
Supplementary information is available for this paper.
Correspondence and requests for materials should be addressed to Tim Althoff (althoff@cs.washington.edu) and Amy X. Zhang (axz@cs.uw.edu).
Supplementary materials
List of supplementary materials
Table LABEL:supp:table:mbfc
Figures S1 to S29
Table S1: Definitions and statistics of factuality and bias categories offered by Media Bias/Fact Check (mediabiasfactcheck.com).
Category Definition # of sources
Factuality:
- Very high: The source is consistently factual, relies on credible information, promptly corrects errors, and has never failed any fact checks in news reporting or opinion pieces. 118
- High: The source is mostly factual and uses mostly credible, low-biased, or high-factual sources. It corrects errors quickly and has failed only one news fact check and up to two op-ed fact checks. 2,313
- Mostly factual: The source is generally accurate but may have a few uncorrected fact-check failures. It can fail up to three op-ed fact checks, especially if it is a low-volume site. While it may use biased sources occasionally, it mostly links to factual content. It is usually pro-science but may sometimes use misleading wording or offer alternative viewpoints. It is reasonably transparent and trustworthy most of the time, but caution is advised. 326
- Mixed: The source may rely on improper sourcing or link to other biased or mixed-factual sources. It often has multiple failed fact checks and does not correct false information or lacks transparency, including the absence of a disclosed mission statement or ownership details. Sources rejecting established scientific consensus on issues such as climate change or vaccines will receive this rating or lower. Sources identified as overt propaganda or designated as hate groups by reputable third-party evaluators will receive this rating or lower due to their inherent bias and potential spread of misleading information. 1,437
- Low: The source is often unreliable and should be fact-checked for fake news, conspiracy theories, and propaganda. 677
- Very low: The source is almost always unreliable and should always be fact-checked for intentional misinformation. 252
Total: 5,123
Bias:
- Least biased: The source has minimal bias and uses very few loaded words (i.e., wording that attempts to influence an audience by using an appeal to emotion or stereotypes). It is factual and usually sourced. 1,054
- Left-center: The source has a slight to moderate liberal bias. It often publishes factual information that utilizes loaded words to favor liberal causes. It is generally trustworthy for information but may require further investigation. 850
- Right-center: Similar to the definition of left-center bias but replacing liberal with conservative. 492
- (Extremely) left: The source is moderately to strongly biased toward liberal causes through story selection or political affiliation. It may utilize strong loaded words, publish misleading reports, and omit reporting of information that may damage liberal causes. It may be untrustworthy. 402
- (Extremely) right: Similar to the definition of (extremely) left bias but replacing liberal with conservative. 314
- Pro-science: The source consists of legitimate science or is evidence-based through the use of credible scientific sourcing. Legitimate science follows the scientific method, is unbiased, and does not use emotional words. The source also respects the consensus of experts in the given scientific field and strives to publish peer-reviewed science. It may have a slight political bias but adheres to scientific principles. 189
- Conspiracy- pseudoscience: The source may publish unverifiable information not always supported by evidence. It may be untrustworthy for credible or verifiable information, so fact-checking and further investigation are recommended on a per-article basis when obtaining information from it. 433
- Questionable: The source exhibits one or more of the following: extreme bias, consistent promotion of propaganda or conspiracies, poor or no sourcing of credible information, a complete lack of transparency, or fake news (i.e., the deliberate attempt to publish hoaxes or disinformation for profit or influence). It may be very untrustworthy and should be fact-checked on a per-article basis. 1,390
- Satire: The source exclusively uses humor, irony, exaggeration, or ridicule to expose and criticize people’s stupidity or vices, particularly in the context of contemporary politics and other topical issues. It does not attempt to deceive. 148
Total: 5,272
Figure S1: Distribution of potential misinformation in X Community Notes (as of February 2023) that received its first response (gray) or first high-quality response (orange) within a certain amount of time.
Figure S2: Examples that show that generating multiple queries helps decompose a post, which may have multiple claims that each needs to be verified, whereas generating one query may overlook some of them and hence lead to not comprehensive identifications of (in)accuracies. Bold text: the verification-needed claims that are overlooked when generating one query but captured when generating more than one query.
Figure S3: Examples that show how retrieved web pages with relatively low relevance to potential misinformation can promote LLM (in this case, GPT-4) hallucinations when generating responses.
Figure S4: Examples of informative image captions, which augment image captions with names of visually represented celebrities and embedded text (see Methods for implementation details).
Figure S5: Examples of Muse-generated responses to accurate, partially accurate, and factually correct but misleading content on social media.
Figure S6: An example of a high-helpfulness response from Community Notes displayed on the corresponding tweet and visible to the public.
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S7: Expert evaluation results for textual posts (#=155).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S8: Expert evaluation results for multimodal (textual and visual) posts (#=77).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S9: Expert evaluation results for political and international-affair posts (#=80).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S10: Expert evaluation results for economic and business posts (#=38).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S11: Expert evaluation results for crime and law posts (#=38).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S12: Expert evaluation results for social-issue and human-right posts (#=30).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S13: Expert evaluation results for health and medicine posts (#=24).
Figure S14: Images as examples used for informative image captioning with in-context learning.
1Describe an image in an informative way. Your description should be only based on the given {short caption}, {name of each person}, and {raw text}. If the image is from social media, you should start with "A screenshot of". If the image is a quote from someone, you should start with "A quote from" followed by this person’s name if there is any, then by the quoted text. If the image is an article, you should start with "An article". If the image is a photo, you should start with "A photo of". If the image is a map, you should start with "A map of". {raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.2short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}3name of each person: {Fran Lebowitz}4raw text: {"In real life, I assure you, there is no such thing as algebra."}5image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}6short caption: {two men in suits}7name of each person: {Jim Caviezel, Michael Emerson}8raw text: {}9image description: {A photo of Jim Caviezel and Michael Emerson in suits}10short caption: {robbie leems on twitter}11name of each person: {}12raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother! 1 2 3,790}13image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}14short caption: {a moose}15name of each person: {}16raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession. finance.yahoo.com Could the stock market power through a recession? ’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}17image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession? ’This would be rare.’" with a picture of a moose. The post was posted at 09:57 22/4/2023.}18short caption: {a person pouring tea into a cup}19name of each person: {}20raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}21image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup. It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}22short caption: {two people standing next to each other with the words love is blind}23name of each person: {Nick Lachey}24raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}25image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other. It was written by Hayley Miller and published by MSNBC.}26short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}27name of each person: {Rahul Kanwal}28raw text: {How engaged are the most-followed journalists on Twitter? Percentage of tweets from each journalist that are at-replies BDUTT 64% RealMikeWilbon 38% 34% ErinAndrews 31% stephenasmith 24% andersoncooper rahulkanwal 13% 13% ninagarcia 5% sardesairajdeep 1% BillSimmons maddow 1% 0% 10% 20% 30% 40% 50% 60% chart by mattmaldre.com}29image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies. The chart was made by mattmaldre.com.}30short caption: {a graph showing the global defense budget by region}31name of each person: {}32raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}33image description: {A graph showing the global defense budget by region. It is from Aviation Week.}34short caption: {[IMAGE_CAPTION]}35name of each person: {[CELEBRITIES]}36raw text: {[OCR]}37image description:
Figure S15: LLM prompt for informative image captioning.
Figure S16: Distribution of helpfulness scores of laypeople’s responses in X Community Notes. All: All laypeople’s responses in Community Notes. Average helpfulness: Laypeople’s responses in Community Notes identified with average helpfulness and used in our study. High helpfulness: Laypeople’s responses in Community Notes identified with high helpfulness and used in our study. For $x\pm y$, $x$: mean, $y$: standard deviation. Community Notes data are regularly updated; ours are up until February 12, 2023.
Figure S17: Distribution of creation times of laypeople’s responses in X Community Notes used in our study. Median of the creation time of laypeople’s average-helpfulness responses: 16 hours after the tweet was posted. Median of the creation time of laypeople’s high-helpfulness responses: 13 hours after the tweet was posted.
Figure S18: Annotation instructions (page 1/7, continued on the next page).
Figure S19: Annotation instructions (page 2/7, continued on the next page).
Figure S20: Annotation instructions (page 3/7, continued on the next page).
Figure S21: Annotation instructions (page 4/7, continued on the next page).
Figure S22: Annotation instructions (page 5/7, continued on the next page).
Figure S23: Annotation instructions (page 6/7, continued on the next page).
Figure S24: Annotation instructions (page 7/7).
Figure S25: Annotation task page.
Figure S26: Impact of starting times of responding to tweets on Muse’s performance. The simulated starting time for Muse (avg): Thirty minutes before the corresponding laypeople’s average-helpfulness responses was created (median: 16 hours after the corresponding tweet was posted; Supplementary Fig. S17). The simulated starting time for Muse (high): Thirty minutes before the corresponding laypeople’s high-helpfulness responses was created (median: 13 hours after the corresponding tweet was posted; Supplementary Fig. S17). The simulated starting time for Muse: The post time of the corresponding tweet (i.e., 0 hours after the corresponding tweet was posted).
Figure S27: Distribution of post times of tweets in X Community Notes included in our analyses.
Figure S28: Impact of post times of tweets on the performance of Muse and baselines.
(a) Overall quality of responses.
(b) Quality of responses in identifying and explaining (in)accuracies.
(c) Quality of responses in generated text.
(d) Quality of responses in references.
[…]