### v4

# suppReferences

Abstract
Real-world misinformation, often multimodal, can be partially correct and factual but misleading by cherry-picking, conflating correlation with causation, and using other tactics. Such misinformation is severely understudied, challenging to address, and harms many social domains, including democracy, economics, justice, human rights, and public health, particularly on social media, where it can spread rapidly. High-quality and timely correction of misinformation that identifies and explains its inaccuracies and accuracies has been shown to effectively reduce false beliefs. Despite the wide acceptance of manual correction, it is difficult to be timely and scalable, a concern as technologies like large language models (LLMs) make misinformation easier to produce. LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce false content, and limitations in addressing multimodal information. We propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information. By retrieving evidence as refutation or supporting context, Muse identifies and explains (in)accuracies in a piece of content—not presupposed to be misinformation—with references. It conducts multimodal retrieval and interprets visual content to verify and correct multimodal content. Given the absence of a systematic and comprehensive evaluation approach, we propose and define 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references. Then, fact-checking experts correspondingly evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts that may or may not be misleading. The results demonstrate Muse’s ability to write high-quality responses to potential misinformation—across modalities, tactics, domains, political leanings, and for information that has not previously been fact-checked online—within minutes of its appearance on social media. Overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%. Our work provides a general methodological and evaluative framework to correct misinformation at scale. It reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
Introduction
Misinformation has far-reaching and detrimental effects on individuals and society[1, 2, 3, 4, 5]. It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[6, 3, 7]. Evidence has demonstrated that election misinformation helped fuel the January 6th attack on the U.S. Capitol, where five people died and more than 100 police officers were injured[8]. Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10]. The result is tragic, where COVID-19 vaccines could have prevented at least 232,000 COVID-19-associated deaths between May 2021 and September 2022 in the U.S. alone[11]. Concerns over misinformation on social media have been particularly significant[12, 13, 14], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[15, 16].
Fortunately, high-quality and timely correction of misinformation, which identifies what part(s) of its content is or is not accurate and explains why that part of the content is (in)accurate with references, has been shown to effectively reduce the spread of misinformation and false beliefs[17, 18, 4, 19]. While domain experts (e.g., on FactCheck.org) and groups of laypeople (e.g., on X Community Notes, formerly Twitter Birdwatch) have played pivotal roles in correcting misinformation[19, 20, 17], keeping pace with massive social media posts is impossible. As Brandolini’s law indicates[21], correcting misinformation is laborious—often requiring domain knowledge, information and media literacy, and explanatory skills—whereas creating and spreading misinformation is easy. As a consequence, 88% of suspicious content on X did not receive any response, and 93% did not receive a high-quality response within the first hour (according to X/Twitter Community Notes as of February 2023; Supp. Fig. S1). Even high-quality responses suffer from limited effectiveness when they are created after rather than before initial bursts of attention to misinformation content[4, 14]. The absence of sufficient coverage also markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[22].
While generative AI models like LLMs raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible. Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand. Indeed, LLMs have revolutionized the field of AI and presented remarkable capabilities across domains and tasks[23, 24, 25]. However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[26]. Specifically, GPT-4[27] (as of March 2023), Llama-3 (as of July 2024)[28], MisinfoCorrect[29], MADR[30], JustiLM[31], and many other LLMs lack access to constantly growing and changing knowledge. These models are thereby ill-equipped to combat misinformation on emerging topics. They either do not provide or “hallucinate” references, which can be fabricated or irrelevant[26, 32]. A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[33, 34, 35, 36, 37, 38, 39]. However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating falsehood and backfiring (i.e., reinforcing rather than reducing false beliefs)[40, 41, 42]. Every above LLM also struggles with counteracting multimodal misinformation due to their nonacceptance of visual inputs.
Real-world misinformation commonly encompasses content that is partially correct and even factual but misleading through various tactics and across domains[2, 3]. Such misinformation can be more difficult to recognize, prevalent, and harmful[2, 3]. Existing LLM and other AI models to tackle misinformation focus narrowly on it that is factually incorrect (e.g., MADR[30], FOLK[38], and ProgramFC[43]), within a single domain (e.g., politics[30] or COVID-19[29, 39]), or employing one specific tactic (e.g., repurposing authentic images with fabricated stories[44]). A general AI framework to practically identify and correct misinformation is demanded but absent. Evaluating the quality of a correction presented in natural language is another core challenge. Related studies heavily rely on automatic evaluation, which calculates the similarity between model-generated and “ground-truth” text using metrics such as ROUGE[31, 44]. However, what qualifies a correction as the ground truth is unclear[29], and existing metrics struggle to capture content factuality, especially when compared to fact-checking experts[26]. The need for an accurate and comprehensive understanding of a model’s performance in identifying and explaining (in)accuracies, generated text, and references is recognized, but a solution is still lacking.
In this article, we propose Muse, a scalable approach for multimodal misinformation correction. Muse makes use of an LLM and augments it with the ability to handle images, access timely and credible knowledge on the web, retrieve evidence that refutes or contextualizes the given content that may or may not be misinformation, and generate clear explanations with accurate and trustworthy references. Muse is a nonparametric approach and therefore easily and cheaply updated, especially compared to parametric language models. Fact-checking experts comprehensively evaluate Muse-generated responses to real social media posts that potentially are misinformation, and compare them to baselines including GPT-4 and high-quality responses based on the collective efforts of laypeople. Our assessment measures the overall quality of a response, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness when it identifies and explains (in)accuracies; the relevance, factuality, fluency, coherence, and toxicity of generated text; and the reachability, relevance, and credibility of references. We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in accurately and promptly responding to potential misinformation. Results demonstrate Muse’s advance no matter whether the potential misinformation content is textual with or without images; fabricates a story, misinterprets or misrepresents facts, lacks context, implies false or oversimplified causation, uses loaded language, presents false or biased data, has improper analogy or equivalence, and adopts other tactics; across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine; leans liberal or conservative; or whether the content has previously fact-checked online or not.
Approach
Muse is designed to automatically respond to content that potentially is misinformation. In other words, the content might be (partially) inaccurate or factually accurate but misleading, all of which are misinformation, or fully accurate as non-misinformation. The content can contain text with or without visuals. The response should identify what part(s) of the content is or is not accurate, explain why that part of the content is or is not accurate, and provide links as references. Overall, Muse features three modules to reliably and efficiently address generalized misinformation: (1) a highly parallelizable relevance- and credibility-aware evidence retriever, (2) an LLM-based evidence-assisted response generator, and (3) an image describer that enables text-only LLMs to jointly interpret multimodal signals from the input.
Fig. 1 illustrates Muse’s pipeline. We start by introducing the details of Muse with a piece of text-only potential misinformation as the input. First, Muse generates queries based on an LLM from the potential misinformation (Fig. 1b; Methods). Each query acts as the input of a web search engine to access timely updated web content and obtain a list of web links directly relevant to the query (Fig. 1b; Methods). After scraping the content from these web links, Muse calculates their direct relevance to the potential misinformation and removes irrelevant web pages (Fig. 1b; Methods). Then, Muse determines the credibility of web pages by looking up their publishers’ factuality and bias ratings and selects pages with high factuality and minimal bias (Fig. 1c; Methods). Next, Muse leverages an LLM to extract text from each of the web pages as evidence. Such evidence can refute the potential misinformation, typically happening when it is misinformation with false claims, or provide additional context, which can demonstrate that the potential misinformation is accurate or part(s) of its claims are accurate (Fig. 1d; Methods). Finally, Muse generates a response to the potential misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig. 1d; Methods).
Note that content, especially content posted on social media, often contains extraneous information that does not need verification and is irrelevant to correction, including unverifiable opinions or emojis, such as the textual content of the false post in Fig. 1. Therefore, generating queries instead of simply using the post content improves web searches as a way of denoising the post content; see example queries in Fig. 1b. Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs verification or correction, whereas generating one query may overlook some of the claims and hence lead to not comprehensive identifications and explanations of (in)accuracies (Supp. Fig. S2). Another concern may arise from filtering retrieved web pages based on how relevant the page content is to the potential misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction. However, the increase in selected web pages drastically elevates the expense of Muse (Methods). We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating responses (Supp. Fig. S3). Moreover, as illustrated in Fig. 1c, Muse filters and ranks the selected web pages by their publishers’ factuality and bias. It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation) or gone through all the credible pages.
When the input is a piece of multimodal (textual and visual) content, Muse first generates textual description of each image (Fig. 1a) so that the content can be handled by any LLM on downstream tasks, including query generation (Fig. 1b), evidence extraction (Fig. 1d), and response generation (Fig. 1d). Specifically, Muse augments image captioning models developed to describe an image in natural language with recognizing celebrities and optical characters based on an LLM (Fig. 1a; Methods). Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with features crucial for making accurate verification and corrections[45] (Methods). For example, even a state-of-the-art image captioning model[46] may describe the visual misinformation in Fig. 1 as simply “list of banned books in Florida.” The description overlooks the listed titles of books that are essential for the visual content’s verification and correction (see more examples in Supp. Fig. S4). In addition, Muse conducts multimodal search on the web and computes multimodal relevance to filter out irrelevant web pages (Fig. 1b; Methods).
Evaluation
Muse’s evaluation was based on X Community Notes data. Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has been shown to reduce the spread of misinformation[17]. Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies. A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supp. Fig. S6). We included the tweets from Community Notes with at least two responses ($n=247$, posted from February 2021 to February 2023); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods). Though we do not presuppose the accuracy of the tweets in Muse’s design and evaluation, we found that more than half of the tweets are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods). We further generated responses to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods). Experts in fact-checking and journalism evaluated the quality of responses by various approaches to the same tweet (Methods); they were blinded to which approach had generated each response. The evaluation contains 13 specific criteria, covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references (Methods). It also contains the overall quality of a response by taking all 13 evaluation criteria into account (Methods).
Our primary finding is that the overall quality of Muse-generated responses is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig. 2a). The overall quality of Muse-generated responses has an average score of 8.1 out of 10, 29% higher than laypeople’s high-helpfulness responses (mean: 6.3; $p=3\times 10^{-48}$, by Mann-Whitney U test unless otherwise specified; $N=464$), 37% higher than GPT-4-generated responses (mean: 5.9; $p=4\times 10^{-42}$; $N=464$), and 56% significantly higher than laypeople’s average-helpfulness responses (mean: 5.2; $p=5\times 10^{-81}$; $N=462$). Despite statistical insignificance between the overall quality of laypeople’s high-helpfulness responses and GPT-4-generated responses ($p=0.4$; $N=464$), the overall quality of GPT-4-generated responses has the highest variability, and GPT-4 generates more responses with extremely low quality. The standard deviation of the overall quality of GPT-4-generated responses is 2.7, vs only 2.0 for Muse and laypeople. 10% of GPT-4’s generated responses have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness responses, 3% for laypeople’s high-helpfulness responses, and 2% for Muse-generated responses. Note that laypeople’s responses were created on average 14 hours after the tweet was posted on social media. Here, Muse only retrieved web pages published before the tweet was posted (Methods).
Examining specific components of response quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in identifying and explaining (in)accuracies (Fig. 2b-f). Experts assessed that Muse-generated responses more explicitly identify and explain where and why a tweet is (in)accurate than GPT-4’s and laypeople’s high-helpfulness responses (Fig. 2b). 89% of Muse’s generated responses explicitly identify and explain (in)accuracies, 16% more than GPT-4-generated responses, 29% more than laypeople’s high-helpfulness responses, and 43% more than laypeople’s average-helpfulness responses (Fig. 2b). As for identifying where a tweet is (in)accurate, we found that Muse more comprehensively identifies a tweet’s (in)accuracies with fewer mistakes—here, mistakes indicate falsely claiming where a tweet should be inaccurate as accurate or where a tweet should be accurate as inaccurate—than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2c-d). 91% of Muse’s generated responses have at least one correct identification without any mistake, 11% more than laypeople’s high-helpfulness responses, 19% more than GPT-4-generated responses, and 26% more than laypeople’s average-helpfulness responses (Fig. 2c). Muse has 61% of generated responses accurately identifying all the (in)accuracies in a tweet, vs GPT-4 has 38%, laypeople who produce high-helpfulness responses have 26%, and laypeople who produce average-helpfulness responses have 17% only (Fig. 2d). Furthermore, Muse explains (in)accuracies more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2e-f). 70% of responses by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness responses, 47% by GPT-4, and 37% for laypeople’s average-helpfulness responses only (Fig. 2e). Meanwhile, the average informativeness score of Muse-generated responses is 7.9, 32% higher than laypeople’s high-helpfulness responses, 36% higher than GPT-4-generated responses, and 65% higher than laypeople’s average-helpfulness responses (Fig. 2f).
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of generated text (Fig. 2g-k). Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=463$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$). Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness responses by laypeople and 45% for GPT-4-generated text (Fig. 2h). Almost all of Muse-generated text does not have any mistake in the use of English (Fig. 2i) and is not biased, impolite, and provoking (Fig. 2k), and 91% is highly coherent and logical (vs 76% and 61% for laypeople, Fig. 2j).
Additionally, results reveal that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of references (Fig. 2l-n). First, GPT-4 hallucinates references frequently. 49% of its links result in “page-not-found” errors (Fig. 2l), and only 76% of reachable links are relevant to the generated text (Fig. 2m). Muse significantly reduces such hallucinations with nearly 100% links being reachable (Fig. 2l) and 96% reachable links being relevant to the generated text (Fig. 2m). Meanwhile, Muse’s references are more credible than the references offered in even high-helpfulness responses by laypeople ($p=4\times 10^{-11}$; $N=744$; Fig. 2n).
We further analyzed the robustness of Muse in addressing diverse forms of real-world misinformation, which may vary in modality, difficulty to correct, political leaning, domain, and adopted tactics. We first observed that the quality of Muse-generated responses to textual and multimodal (textual and visual) content is consistently higher than responses by GPT-4 and even high-helpfulness responses by laypeople, with a margin of 21% or higher (Fig. 3a). GPT-4 can be comparable to laypeople who write high-helpfulness responses when responding to text-only tweets. However, its performance significantly declines when responding to multimodal tweets, often resulting in a quality lower than laypeople who produce average-helpfulness responses.
Second, we removed tweets where Muse’s response references at least one related fact-checking article, accounting for 16% of all tweets. Notably, in only 6% of all tweets, Muse’s response exclusively references related fact-checking article(s). The quality of Muse-generated responses to these tweets can be more associated with Muse’s ability to summarize retrieved professional fact-checking articles, rather than fact-checking such content from scratch. We found that the quality of Muse-generated responses to content that has not been fact-checked online is at least 28% higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig. 3b). The average quality score of Muse-generated responses for these potentially more challenging tweets is 8.0, close to its 8.1 average score for all tweets.
Third, LLMs have exhibited political bias [47], which may diminish their ability to debunk misinformation that aligns with their political stance. We annotated the political leaning of tweets and found that GPT-4, which has an observed liberal bias [47], indeed performs 5% worse in correcting misinformation that aligns with liberal ideology compared to conservative ideology (Fig. 3c). Nevertheless, the quality of Muse-generated responses to both liberal and conservative content has the same average score of 8.3. This score is consistently higher than GPT-4-generated responses and even high-helpfulness responses from laypeople by at least 26%.
Finally, we annotated the domain of tweets and tactics used to make them or part of them false or misleading. Results in Fig. 3d-e indicate that the quality of Muse-generated responses to potential misinformation across different domains and tactics maintains an average score around 8, consistently higher than responses by GPT-4 and even high-helpfulness responses from laypeople, with a margin of at least 19%. Here, domains broadly include politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine. Tactics primarily include fabricating, e.g., a story with or without an authentic image, or using a digitally fabricated or altered screenshot or photo; misinterpreting or misrepresenting someone’s claim, a symbol, a policy, and others; lacking context; implying false or oversimplified causation, such as conflating correlation with causation; using loaded language; presenting false, partial, or biased data, such as cherry-picking; and having improper analogy or equivalence (see examples in Supp. Fig. S7).
Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[48, 49], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[2, 3]. Correcting misinformation requires identifying what part(s) of the content is (in)accurate and explaining why that part of the content (in)accurate with trustworthy references. Such identifications with clear explanations have been shown to reduce misinformation’s spread and people’s false beliefs, but its effectiveness is affected by quality, timeliness, and scalability[4, 22]. However, research on scaling up and accelerating misinformation correction is still at the early stage. Existing AI models are ill-equipped to identify and correct misinformation on social media. Such misinformation often spreads rapidly and is not restricted to narrow domains. It can combine (in)accurate, factually accurate but misleading, and unverifiable claims. It can employ various tactics, deliberately or accidentally, to create falsehood or misleadingness. Addressing it requires comprehending content that can be multimodal and the context beyond it that often involves emerging events. Existing methods to evaluate the quality of corrections presented in natural language also struggle to be accurate and comprehensive. Such evaluation relies on professional knowledge and skills. It should thoroughly assess the identification and explanation of (in)accuracies, as well as the use of references from diverse perspectives.
We propose Muse, which augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references, and 13 dimensions of misinformation correction quality. Evaluation by fact-checking experts demonstrates the high quality of Muse’s automatically generated responses to social media posts that potentially are misinformation. Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining (in)accuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4. Note that addressing LLM ’hallucination’ has been a long-standing challenge stressed in many works[26, 50].
We provide Muse as a solution to assist social media users and platforms in reliably, scalably, promptly, and transparently responding to suspicious content. Muse is end-to-end and thereby simple to use. It is nonparametric and thereby easily and cheaply updated[51]. Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities, domains, political leanings, and tactics and that has not even been fact-checked online (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated responses show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating responses with the least toxic text and most credible references[29, 41, 42] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[22] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. These high-quality responses by Muse can generally be obtained within minutes of suspicious content appearing on social media (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f). Muse can also inform users of the current lack of evidence, provide an accuracy nudge, and when applicable, express uncertainty (Supp. Fig. S8).
Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (as of February 2024; Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the task’s complexity and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[19] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation. Notably, we also changed Muse’s foundation LLM from GPT-4 to open-sourced Llama-3 (70B) and observed two Muses are comparable without modifying any other implementation details (Supp. Fig. S9; Methods), suggesting Muse’s generalizability. Note that the knowledge cutoff for Llama-3 is December 2023, whereas that of GPT-4 is September 2021. Llama-3’s training data might have included some or even all tweets used in evaluation and related events, knowledge, and corrections (e.g., their Community Notes data). As a comparison, GPT-4’s training data might only include up to 10% (Methods, where we show that Muse’s performance remains consistent after removing these tweets). Hence, it is unfair to compare between Llama-3-based and GPT-4-based Muse, and we did not include Llama-3 and its based Muse in expert evaluation. Nevertheless, our qualitative analysis illustrated in Supp. Fig. S9 reveals the potential to significantly reduce Muse’s cost, as GPT-4 was Muse’s most costly component. Furthermore, as one of the most capable open-sourced LLMs, Llama-3 still cannot accept visual inputs (as of August 2024), emphasizing the importance of Muse’s image describer module.
This work also faces the following limitations. First, although Muse is capable of responding to multimodal misinformation with text and images, it cannot accept video inputs. Second, we only focus on English, one of the most spoken languages in the world. Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has been shown to reduce the spread of misinformation[17] and transparent. X is also a popular social media platform, and one where more than half of users consume news regularly[52] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one foundation LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (see Methods and Supp. Fig. S24 for additional results of our ablation study). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[53, 54].
Conclusion
We proposed Muse, a nonparametric LLM to identify and explain (in)accuracies in social media content. Muse can integrate textual and visual information, access timely knowledge, and generate responses that use natural languages and have accurate and trustworthy references. Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who write high-helpfulness responses on X Community Notes in identifying and explaining (in)accuracies, generating high-quality text, providing high-quality references, and producing high-quality responses overall. Muse excels when the responded content is textual or multimodal, related to a broad range of domains, false or misleading by applying various tactics, and even when the content has not been fact-checked online. Muse also performs consistently well when responding to liberal and conservative content. This work demonstrates the potential of LLMs to respond to online misinformation in a scalable, prompt, reliable, and transparent manner—addressing key bottlenecks where existing studies struggle. This work also contributes to a comprehensive framework, incorporating with 13 associated standards (Methods, where we analyzed their importance), for evaluating misinformation correction models. Additionally, we publicly release rich annotations on social media posts covering aspects such as political leanings, domains, and tactics that make them false or misleading, along with expert-annotated responses to enable research.
Acknowledgements
We would like to thank Hacks/Hackers for advertising the study and helping with the recruitment. We also thank members of the UW Social Futures Lab and Behavioral Data Science Group for their suggestions and feedback. This work was supported in part by the National Science Foundation’s Convergence Accelerator program under Award No. 49100421C0037. T.A. and A.S. were supported in part by the Office of Naval Research (#N00014-21-1-2154), NSF IIS-1901386, and NSF CAREER IIS-2142794.
Author contributions
X.Z., A.X.Z., and T.A. designed the study. X.Z. led, and A.S. assisted in developing the model. A.S. built the web interface for evaluation. X.Z. prepared and analyzed the data. X.Z. drafted and all authors revised the manuscript. A.X.Z. and T.A. supervised the study.
Competing interests
The authors declare no competing interests.
Additional information
Supplementary information is available for this paper.
Correspondence and requests for materials should be addressed to Tim Althoff (althoff@cs.washington.edu) and Amy X. Zhang (axz@cs.uw.edu).
Supplementary materials
List of supplementary materials
Table LABEL:supp:table:mbfc
Figures S1 to S25
Table S1: Definitions and statistics of factuality and bias categories offered by Media Bias/Fact Check (mediabiasfactcheck.com).
Category Definition # of sources
Factuality:
- Very high: The source is consistently factual, relies on credible information, promptly corrects errors, and has never failed any fact checks in news reporting or opinion pieces. 118
- High: The source is mostly factual and uses mostly credible, low-biased, or high-factual sources. It corrects errors quickly and has failed only one news fact check and up to two op-ed fact checks. 2,313
- Mostly factual: The source is generally accurate but may have a few uncorrected fact-check failures. It can fail up to three op-ed fact checks, especially if it is a low-volume site. While it may use biased sources occasionally, it mostly links to factual content. It is usually pro-science but may sometimes use misleading wording or offer alternative viewpoints. It is reasonably transparent and trustworthy most of the time, but caution is advised. 326
- Mixed: The source may rely on improper sourcing or link to other biased or mixed-factual sources. It often has multiple failed fact checks and does not correct false information or lacks transparency, including the absence of a disclosed mission statement or ownership details. Sources rejecting established scientific consensus on issues such as climate change or vaccines will receive this rating or lower. Sources identified as overt propaganda or designated as hate groups by reputable third-party evaluators will receive this rating or lower due to their inherent bias and potential spread of misleading information. 1,437
- Low: The source is often unreliable and should be fact-checked for fake news, conspiracy theories, and propaganda. 677
- Very low: The source is almost always unreliable and should always be fact-checked for intentional misinformation. 252
Total: 5,123
Bias:
- Least biased: The source has minimal bias and uses very few loaded words (i.e., wording that attempts to influence an audience by using an appeal to emotion or stereotypes). It is factual and usually sourced. 1,054
- Left-center: The source has a slight to moderate liberal bias. It often publishes factual information that utilizes loaded words to favor liberal causes. It is generally trustworthy for information but may require further investigation. 850
- Right-center: Similar to the definition of left-center bias but replacing liberal with conservative. 492
- (Extremely) left: The source is moderately to strongly biased toward liberal causes through story selection or political affiliation. It may utilize strong loaded words, publish misleading reports, and omit reporting of information that may damage liberal causes. It may be untrustworthy. 402
- (Extremely) right: Similar to the definition of (extremely) left bias but replacing liberal with conservative. 314
- Pro-science: The source consists of legitimate science or is evidence-based through the use of credible scientific sourcing. Legitimate science follows the scientific method, is unbiased, and does not use emotional words. The source also respects the consensus of experts in the given scientific field and strives to publish peer-reviewed science. It may have a slight political bias but adheres to scientific principles. 189
- Conspiracy- pseudoscience: The source may publish unverifiable information not always supported by evidence. It may be untrustworthy for credible or verifiable information, so fact-checking and further investigation are recommended on a per-article basis when obtaining information from it. 433
- Questionable: The source exhibits one or more of the following: extreme bias, consistent promotion of propaganda or conspiracies, poor or no sourcing of credible information, a complete lack of transparency, or fake news (i.e., the deliberate attempt to publish hoaxes or disinformation for profit or influence). It may be very untrustworthy and should be fact-checked on a per-article basis. 1,390
- Satire: The source exclusively uses humor, irony, exaggeration, or ridicule to expose and criticize people’s stupidity or vices, particularly in the context of contemporary politics and other topical issues. It does not attempt to deceive. 148
Total: 5,272
Figure S1: Distribution of potential misinformation in X Community Notes (as of February 2023) that received its first response (gray) or first high-quality response (orange) within a certain amount of time.
Figure S2: Examples that show that generating multiple queries helps decompose a post, which may have multiple claims that each needs to be verified, whereas generating one query may overlook some of them and hence lead to not comprehensive identifications of (in)accuracies. Bold text: the verification-needed claims that are overlooked when generating one query but captured when generating more than one query.
Figure S3: Examples that show how retrieved web pages with relatively low relevance to potential misinformation can promote LLM (in this case, GPT-4) hallucinations when generating responses.
Figure S4: Examples of informative image captions, which augment image captions with names of visually represented celebrities and embedded text (see Methods for implementation details).
Figure S5: Examples of Muse-generated responses to accurate, partially accurate, and factually correct but misleading content on social media.
Figure S6: An example of a high-helpfulness response from Community Notes displayed on the corresponding tweet and visible to the public.
Figure S7: Examples of social media content that uses various tactics to make it or part of it false or misleading. These tactics include a: Fabricating a news event (#1), a story with an authentic image (#2, also related to conspiracy theory), and an image with AI (#3). b: Lacking context regarding the motivation behind behavior. Here, it is the consideration that the bill contained no funding for Florida, without which could mislead readers. c: Misinterpreting or misrepresenting a plan, where Biden did not promise to cancel all the debt (#1), or symbol, which is not Nazi SS runes but a shorthand for the 46th Separate Airmobile Brigade (#2). d: Using loaded language. e: Improperly analogizing or equating masks blocking respiratory droplets and underwear blocking gas molecules (#1), and COVID-19 and cancer (#2). f: Presenting false, partial, or biased data. 365348 includes multiple reports for the same child and runaways who may not be considered missing because of their guardians know their whereabouts (#1). The data ends in 1885 and is from a specific high elevation site in Greenland, not global temperatures (#2). g: Implying false or oversimplified causation. Many COVID-19 deaths occurred among Medicare beneficiaries because Medicare primarily serves the groups who are at greater risk of adverse outcomes from COVID-19 (#1). The boy was shot when—but not because—he was campaigning for Raphael Warnock (#2). Note that one content may apply more than one tactic intentionally or accidentally.
Figure S8: Examples of Muse-generated responses to social media posts that lack evidence to verify.
Figure S9: Comparison between responses generated by GPT-4, Llama-3, and their based Muse. We observed that GPT4-based Muse and Llama3-based Muse frequently generate similar responses to social media content (see the top example). GPT4-based Muse occasionally generates more comprehensive corrections than Llama3-based Muse (see the middle example). Note that the knowledge cutoff for Llama3 is December 2023, whereas that of GPT-4 is September 2021. In other words, Llama-3’s training data include more recent events and knowledge than GPT-4’s (see the bottom example, where the post was posted in February 2023, and Llama-3 more accurately identifies its falsehood than GPT-4).
Figure S10: Images as examples used for informative image captioning with in-context learning.
1Describe an image in an informative way. Your description should be only based on the given {short caption}, {name of each person}, and {raw text}. If the image is from social media, you should start with "A screenshot of". If the image is a quote from someone, you should start with "A quote from" followed by this person’s name if there is any, then by the quoted text. If the image is an article, you should start with "An article". If the image is a photo, you should start with "A photo of". If the image is a map, you should start with "A map of". {raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.2short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}3name of each person: {Fran Lebowitz}4raw text: {"In real life, I assure you, there is no such thing as algebra."}5image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}6short caption: {two men in suits}7name of each person: {Jim Caviezel, Michael Emerson}8raw text: {}9image description: {A photo of Jim Caviezel and Michael Emerson in suits}10short caption: {robbie leems on twitter}11name of each person: {}12raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother! 1 2 3,790}13image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}14short caption: {a moose}15name of each person: {}16raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession. finance.yahoo.com Could the stock market power through a recession? ’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}17image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession? ’This would be rare.’" with a picture of a moose. The post was posted at 09:57 22/4/2023.}18short caption: {a person pouring tea into a cup}19name of each person: {}20raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}21image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup. It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}22short caption: {two people standing next to each other with the words love is blind}23name of each person: {Nick Lachey}24raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}25image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other. It was written by Hayley Miller and published by MSNBC.}26short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}27name of each person: {Rahul Kanwal}28raw text: {How engaged are the most-followed journalists on Twitter? Percentage of tweets from each journalist that are at-replies BDUTT 64%29image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies. The chart was made by mattmaldre.com.}30short caption: {a graph showing the global defense budget by region}31name of each person: {}32raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}33image description: {A graph showing the global defense budget by region. It is from Aviation Week.}34short caption: {[IMAGE_CAPTION]}35name of each person: {[CELEBRITIES]}36raw text: {[OCR]}37image description:
Figure S11: LLM prompt for informative image captioning.
Figure S12: Distribution of helpfulness scores of laypeople’s responses in X Community Notes. All: All laypeople’s responses in Community Notes. Average helpfulness: Laypeople’s responses in Community Notes identified with average helpfulness and used in our study. High helpfulness: Laypeople’s responses in Community Notes identified with high helpfulness and used in our study. For $x\pm y$, $x$: mean, $y$: standard deviation. Community Notes data are regularly updated; ours are up until February 12, 2023.
Figure S13: Distribution of creation times of laypeople’s responses in X Community Notes used in our study. Median of the creation time of laypeople’s average-helpfulness responses: 16 hours after the tweet was posted. Median of the creation time of laypeople’s high-helpfulness responses: 13 hours after the tweet was posted.
Figure S14: Annotation instructions (page 1/7, continued on the next page).
Figure S15: Annotation instructions (page 2/7, continued on the next page).
Figure S16: Annotation instructions (page 3/7, continued on the next page).
Figure S17: Annotation instructions (page 4/7, continued on the next page).
Figure S18: Annotation instructions (page 5/7, continued on the next page).
Figure S19: Annotation instructions (page 6/7, continued on the next page).
Figure S20: Annotation instructions (page 7/7).
Figure S21: Annotation task page.
Figure S22: Impact of starting times of responding to tweets on Muse’s performance. The simulated starting time for Muse (avg): Thirty minutes before the corresponding laypeople’s average-helpfulness responses was created (median: 16 hours after the corresponding tweet was posted; Supplementary Fig. S13). The simulated starting time for Muse (high): Thirty minutes before the corresponding laypeople’s high-helpfulness responses was created (median: 13 hours after the corresponding tweet was posted; Supplementary Fig. S13). The simulated starting time for Muse: The post time of the corresponding tweet (i.e., 0 hours after the corresponding tweet was posted).
Figure S23: Impact of post times of tweets on the performance of Muse and baselines.
(a) Overall quality of responses.
(b) Quality of responses in identifying and explaining (in)accuracies.
(c) Quality of responses in generated text.
(d) Quality of responses in references.
[…]
(a) Correlation for all responses.
(b) Correlation for responses whose overall quality is relatively low vs high.
Figure S25: Spearman correlation coefficient between overall response quality and each criterion that specifies overall response quality.