### v5

# Correcting misinformation on social media with a large language model

Abstract
Real-world information, often multimodal, can be misinformed or potentially misleading due to factual errors, outdated claims, missing context, misinterpretation, and more. Such “misinformation” is understudied, challenging to address, and harms many social domains—particularly on social media, where it can spread rapidly. Manual correction that identifies and explains its (in)accuracies is widely accepted but difficult to scale. While large language models (LLMs) can generate human-like language that could accelerate misinformation correction, they struggle with outdated information, hallucinations, and limited multimodal capabilities. We propose Muse, an LLM augmented with vision-language modeling and web retrieval over relevant, credible sources to generate responses that determine whether and which part(s) of the given content can be misinformed or potentially misleading, and to explain why with grounded references. We further define a comprehensive set of rubrics to measure response quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references. Results show that Muse consistently produces high-quality outputs across diverse social media content (e.g., modalities, domains, political leanings), including content that has not previously been fact-checked online. Overall, Muse outperforms GPT-4 by 37% and even high-quality responses from social media users by 29%. Our work provides a general methodological and evaluative framework for correcting misinformation at scale.
Keywords
Misinformation, large language model, retrieval-augmented generation, multimodality, social media
Significance Statement
Misinformed and potentially misleading content are pervasive on social media but difficult to address at scale due to multimodality, diversity, and rapidly emerging topics. Such content can undermine democratic processes, economic stability, public health, and national security. We introduce a large language model augmented with vision-language modeling and credibility-aware web retrieval, along with a comprehensive evaluation framework. The model identifies and explains potentially misleading claims with grounded references, enabling transparent corrections that people can understand and audit. We demonstrate effectiveness and generalizability on real-world social media content through evaluations with fact-checking experts and user perception studies.
Introduction
Misinformation has far-reaching and detrimental effects on individuals and society[1, 11, 3, 4, 5, 6]. It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[7, 3, 8]. Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10]. Concerns over misinformation on social media have been particularly significant[11, 12, 13], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[14, 15].
Fortunately, high-quality and timely correction of misinformation—identifying what part(s) of its content is or is not accurate and explaining why, with supporting references, has been shown to effectively reduce the spread of misinformation and false beliefs[16, 17, 4, 18]. While domain experts (e.g., on FactCheck.org) and groups of laypeople (e.g., on X Community Notes, formerly Twitter Birdwatch) have played pivotal roles in correcting misinformation[18, 19, 16], keeping pace with massive social media posts is impossible. As Brandolini’s law indicates[20], correcting misinformation is laborious—often requiring domain knowledge, information and media literacy, and explanatory skills—whereas creating and spreading misinformation is easy. As a consequence, 88% of suspicious content on X did not receive any response, and 93% did not receive a high-quality response within the first hour (according to X/Twitter Community Notes as of February 2023; Supp. Fig. S1). Even high-quality responses suffer from limited effectiveness when they are created after rather than before initial bursts of attention to misinformation content[4, 13]. The absence of sufficient coverage also markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[21].
While generative artificial intelligence (AI) models like large language models (LLMs) raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible. Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand. However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[22]. Specifically, GPT-4[23] (as of March 2023), Llama-3 (as of July 2024)[24], MisinfoCorrect[25], MADR[26], JustiLM[27], and many other LLMs lack access to constantly growing and changing knowledge. These models are thereby ill-equipped to combat misinformation on emerging topics. They either do not provide or “hallucinate” references, which can be fabricated or irrelevant[22, 28]. A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[29, 30, 31, 32, 33, 34, 35]. However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating falsehood and backfiring (i.e., reinforcing rather than reducing false beliefs)[36, 37, 38]. Every above LLM also struggles with counteracting multimodal misinformation due to their nonacceptance of visual inputs.
Real-world misinformation commonly encompasses content that is partially correct and even factual but misleading through various tactics (e.g., cherry-picking, conflating correlation with causation) and across domains[11, 3]. Such misinformation can be more difficult to recognize, prevalent, and harmful[11, 3]. Existing AI models to tackle misinformation focus narrowly on it that is factually incorrect (e.g., SAFE[39], MADR[26], FOLK[34], ProgramFC[40]), within a single domain (e.g., politics[26] or COVID-19[25, 35]), or employing one specific tactic (e.g., repurposing authentic images with fabricated stories[41]). A general AI framework to practically identify and correct misinformation is demanded but absent. Evaluating the quality of a correction in natural language is another core challenge. Related studies heavily rely on automatic evaluation, which calculates the similarity between model-generated and “ground-truth” text using metrics like ROUGE[27, 41]. However, what qualifies a correction as the ground truth is unclear[25], and existing metrics struggle to capture content factuality, especially when compared to fact-checking experts[22]. Furthermore, how the general public perceives and reacts to AI-generated corrections remains unmeasured. The need for an accurate and comprehensive understanding of a model’s performance in identifying and explaining (in)accuracies, generated text, and references is recognized, but a solution is still lacking.
In this article, we propose Muse, a scalable approach for multimodal misinformation correction. Muse makes use of an LLM and augments it with the ability to handle images, access timely and credible knowledge on the web, retrieve evidence that refutes or contextualizes the given content that may or may not be misinformation, and generate clear explanations with accurate and trustworthy references. Muse does not need to finetune any AI models and therefore easily and cheaply updated. Fact-checking experts comprehensively evaluate Muse-generated responses to real social media posts that potentially are misinformation, and compare them to baselines including GPT-4 and high-quality responses based on the collective efforts of laypeople. Our assessment measures the overall quality of a response, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness when it identifies and explains (in)accuracies; the relevance, factuality, fluency, coherence, and toxicity of generated text; and the reachability, relevance, and credibility of references. We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in accurately and promptly responding to potential misinformation, exhibiting high generalizability. A study with 988 participants measuring end-user perceptions further demonstrates Muse’s superior performance in increasing people’s ability to correctly identify misinformation by 9.8%.
Approach
Muse is designed to automatically respond to content that potentially is misinformation. In other words, the content might have multiple inaccuracies, be (partially) inaccurate, or be factual but potentially misleading, all of which are misinformation, or fully accurate as non-misinformation. The content can contain text with or without visuals. The response should identify which part(s) of the content is or is not accurate, explain why, and provide reference links. Overall, Muse consists of three interconnected modules to address misinformation in real-world settings: (1) an LLM for response generation, (2) a hierarchical, credibility-aware web retriever for evidence extraction, and (3) a multimodal integrator that equips modules 1&2 with multimodal capabilities. Fig. 1 illustrates Muse’s pipeline.
We start by introducing the details of Muse with a piece of text-only potential misinformation as the input. First, Muse generates queries from the potential misinformation (Fig. 1b; Methods). Each query acts as the input of a web search engine to access timely updated web content and obtain a list of web links directly relevant to the query (Fig. 1b; Methods). After scraping the content from these web links, Muse calculates their direct relevance to the potential misinformation and removes irrelevant web pages (Fig. 1b; Methods). Then, Muse determines the credibility of web pages by looking up their publishers’ factuality and bias ratings and selects pages with high factuality and minimal bias (Fig. 1c; Methods). Next, Muse extracts text from each of the web pages as evidence. Such evidence can refute the potential misinformation, typically happening when it is misinformation with false claims, or provide additional context, which can demonstrate that the potential misinformation is accurate or part(s) of its claims are accurate (Fig. 1d; Methods). Finally, Muse generates a response to the potential misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig. 1d; Methods).
Note that content, especially on social media, often contains extraneous information that does not need verification and is irrelevant to correction, including unverifiable opinions or emojis, such as the false post in Fig. 1. Therefore, generating queries instead of simply using the post content improves web searches as a way of denoising the post content; see example queries in Fig. 1b. Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs verification or correction, whereas generating one query may overlook some of the claims and hence lead to not comprehensive identifications and explanations of (in)accuracies (Supp. Fig. S2). Another concern may arise from filtering retrieved web pages based on how relevant the page content is to the potential misinformation: hypothetically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction. However, the increase in selected web pages drastically elevates the expense of Muse (Methods). We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating responses (Supp. Fig. S3). Moreover, as illustrated in Fig. 1c, Muse filters and ranks the selected web pages by their publishers’ factuality and bias. It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation) or gone through all the credible pages.
When the input is a piece of textual and visual content, Muse generates textual description of each image (Fig. 1a) so that the content can be handled by any LLM on downstream tasks, such as response generation. Specifically, Muse augments image captioning models with features that existing models often fail to capture[42] but are crucial for accurate verification and correction (e.g., recognition of celebrities and optical characters; Fig.1a; Methods). For example, even a state-of-the-art image captioning model[43] may describe the visual misinformation in Fig. 1 as simply “list of banned books in Florida.” The description overlooks the listed titles of books that are essential for the visual content’s verification and correction (see more examples in Supp. Fig. S4). In addition, Muse conducts multimodal search on the web and computes multimodal relevance to filter out irrelevant web pages (Fig. 1b; Methods).
Evaluation
Musewas evaluated on X Community Notes data. Community Notes empowers social media users, often laypeople, to collaboratively fact-check posts. It has been shown to reduce the spread of misinformation[16] and is becoming increasingly impactful as it expands to other platforms, such as Meta’s[12]. Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies. A response with a sufficiently high helpfulness score is then displayed on the corresponding post and publicly visible[16] (Supp. Fig. S6). We included the posts from Community Notes with at least two responses ($n=247$, posted from February 2021 to February 2023); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods). Though we do not presuppose the accuracy of the posts in Muse’s design and evaluation, we found that more than half of the posts are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods). We further generated responses to these posts based on Muse and Muse’s variants, including GPT-4 (as of June 2023) (Methods). Next, we defined 13 specific criteria—covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references—to comprehensively assess response quality (Methods). Given the professional knowledge required for such evaluations, we recruited experts in fact-checking and journalism to assess and compare responses generated by various approaches to the same post, using each criterion and an overall quality score that considers all 13 criteria (Methods). The evaluators were blinded to which approach generated each response.
Our primary finding is that the overall quality of Muse-generated responses is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig. 2a). The overall quality of Muse-generated responses has an average score of 8.1 out of 10, 29% higher than laypeople’s high-helpfulness responses (mean: 6.3; $p=3\times 10^{-48}$, by Mann-Whitney U test unless otherwise specified; $N=464$), 37% higher than GPT-4-generated responses (mean: 5.9; $p=4\times 10^{-42}$; $N=464$), and 56% significantly higher than laypeople’s average-helpfulness responses (mean: 5.2; $p=5\times 10^{-81}$; $N=462$). Despite statistical insignificance between the overall quality of responses from laypeople (high-helpfulness) and GPT-4 ($p=0.4$; $N=464$), the overall quality of GPT-4’s responses has the highest variability, and GPT-4 generates more responses with extremely low quality. The standard deviation of the overall quality of GPT-4 responses is 2.7, vs only 2.0 for Muse and laypeople. 10% of GPT-4’s responses have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness responses, 3% for laypeople’s high-helpfulness responses, and 2% for Muse. Note that laypeople’s responses were created on average 14 hours after the post was posted on social media. Here, Muse only retrieved web pages published before the post was posted (Methods).
Examining specific components of response quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in identifying and explaining (in)accuracies (Fig. 2b-f). 89% of Muse’s responses explicitly identify and explain (in)accuracies, 16% more than GPT-4’s, 29% more than laypeople’s high-helpfulness responses, and 43% more than laypeople’s average-helpfulness responses (Fig. 2b). As for identifying where a post is (in)accurate, we found that Muse more comprehensively identifies a post’s (in)accuracies with fewer mistakes—here, mistakes indicate falsely claiming where a post should be inaccurate as accurate or where a post should be accurate as inaccurate—than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2c-d). 91% of Muse’s responses have at least one correct identification without any mistake, 11% more than laypeople’s high-helpfulness responses, 19% more than GPT-4, and 26% more than laypeople’s average-helpfulness responses (Fig. 2c). Muse has 61% of responses accurately identifying all the (in)accuracies in a post, vs GPT-4 has 38%, laypeople (high helpfulness) have 26%, and laypeople (average helpfulness) have 17% only (Fig. 2d). Furthermore, Muse explains (in)accuracies more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2e-f). 70% of responses by Muse have fully accurate explanations, vs 55% for laypeople (high helpfulness), 47% by GPT-4, and 37% for laypeople (average helpfulness) only (Fig. 2e). Meanwhile, the average informativeness score of Muse’s responses is 7.9, 32% higher than laypeople (high helpfulness), 36% higher than GPT-4, and 65% higher than laypeople (average helpfulness) (Fig. 2f).
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of generated text (Fig. 2g-k). Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded post ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=463$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$). Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness responses by laypeople and 45% for GPT-4-generated text (Fig. 2h). Almost all of Muse-generated text does not have any mistake in the use of English (Fig. 2i) and is not biased, impolite, and provoking (Fig. 2k), and 91% is highly coherent and logical (vs 76% and 61% for laypeople, Fig. 2j).
Additionally, results reveal that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of references (Fig. 2l-n). First, GPT-4 hallucinates references frequently. 49% of its links result in “page-not-found” errors (Fig. 2l), and only 76% of reachable links are relevant to the generated text (Fig. 2m). Muse significantly reduces such hallucinations with nearly 100% links being reachable (Fig. 2l) and 96% reachable links being relevant to the generated text (Fig. 2m). Meanwhile, Muse’s references are more credible than the references offered in even high-helpfulness responses by laypeople ($p=4\times 10^{-11}$; $N=744$; Fig. 2n).
We further analyzed the generalizability of Muse in addressing diverse forms of real-world misinformation, which may vary in modality, difficulty to correct, political leaning, domain, and adopted tactics. We first observed that the quality of Muse-generated responses to textual and multimodal (+visual) content is consistently higher than GPT-4 and even high-helpfulness responses by laypeople, with a margin of 21% or higher (Fig. 3a). GPT-4 can be comparable to laypeople (high helpfulness) when responding to text-only posts. However, its performance significantly declines when responding to multimodal posts, often resulting in a quality lower than laypeople (average helpfulness).
Second, we removed posts where Muse’s response references at least one related fact-checking article, accounting for 16% of all posts. Notably, in only 6% of all posts, Muse’s response exclusively references related fact-checking article(s). The quality of Muse-generated responses to these posts can be more associated with Muse’s ability to summarize retrieved professional fact-checking articles, rather than fact-checking such content from scratch. We found that the quality of Muse-generated responses to content that has not been fact-checked online is at least 28% higher than GPT-4 and even high-helpfulness responses by laypeople (Fig. 3b). The average quality score of Muse’s responses for these potentially more challenging posts is 8.0, close to its 8.1 average score for all posts.
Third, LLMs have exhibited political bias [45], which may diminish their ability to debunk misinformation that aligns with their political stance. We annotated the political leaning of posts and found that GPT-4, which has an observed liberal bias [45], indeed performs 5% worse in correcting misinformation that aligns with liberal ideology compared to conservative ideology (Fig. 3c). Nevertheless, the quality of Muse-generated responses to both liberal and conservative content has the same average score of 8.3. This score is consistently higher than GPT-4 and even high-helpfulness responses from laypeople by at least 26%.
Finally, we annotated the domain of posts and tactics used to make them or part of them false or misleading. Results in Fig. 3d-e indicate that the quality of Muse’ responses to potential misinformation across different domains and tactics maintains an average score around 8, consistently higher than responses by GPT-4 and even high-helpfulness responses from laypeople, with a margin of at least 19%. Here, domains broadly include politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine. Tactics primarily include fabricating, e.g., a story with or without an authentic image, or using a digitally fabricated or altered screenshot or photo; misinterpreting or misrepresenting someone’s claim, a symbol, a policy, and others; lacking context; implying false or oversimplified causation, such as conflating correlation with causation; using loaded language; presenting false, partial, or biased data, such as cherry-picking; and having improper analogy or equivalence (see examples in Supp. Fig. S7).
We further investigated the impact of potential training data contamination and poster popularity, demonstrating that Muse’s superior performance generalizes to correcting misinformation not included in its training data, as well as misinformation posted by users with varying levels of popularity (Methods). We also conducted a detailed ablation study, showing the effectiveness of the proposed retrieval mechanism and multimodal integration (Methods). Beyond expert evaluation, we conducted an additional study with 988 laypeople to examine how social media users perceive and react to corrections generated by Muse (Methods). The results demonstrate Muse’s superior performance in generating corrections useful to end users. Notably, it is the only approach shown to significantly increase people’s capability to correctly identify misinformation, with an improvement of 9.8%, compared to even strong baselines, including GPT-4 and high-quality responses from laypeople.
Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[46, 47], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[11, 3]. Correcting misinformation requires identifying which part(s) of the content are (in)accurate and explaining why with trustworthy references. This approach can reduce misinformation sharing and false beliefs, but effectiveness depends on the correction’s quality, timeliness, and scalability [4, 21]. Existing approaches, still at an early stage, struggle in particular with complex, nuanced misinformation that spans domains and goes beyond straightforward factual errors, especially when addressing it requires comprehending multimodal content and situating it in rapidly evolving real-world context. Existing methods for evaluating the quality of natural-language corrections also struggle to be accurate and comprehensive. Such evaluation relies on rigorous assessment of the identification and explanation of (in)accuracies, as well as the use of references from diverse perspectives. Beyond content validity, evaluation should also probe likely real-world impact by examining how end users perceive and act on the correction.
We propose Muse, which augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references, as well as 13 dimensions of correction quality. Evaluation by fact-checking experts demonstrates the high quality of Muse-generated responses to social media posts as potential misinformation. Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining (in)accuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4. Note that addressing LLM ’hallucination’ has been a long-standing challenge stressed in many works[22, 48].
As more and more social media platforms (e.g., Facebook, Instagram, Threads [12]) are adopting Community Notes-like approaches to address misinformation, we provide Muse as a potential solution to assist social media users in drafting responses and to help platforms directly respond to misinformation reliably, promptly, and transparently. Muse is end-to-end and thereby simple to use. It is nonparametric and thereby easily and cheaply updated[49]. Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities, domains, political leanings, and tactics and that has not even been fact-checked online (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated responses show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating responses with the least toxic text and most credible references[25, 37, 38] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[21] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. Additionally, we empirically demonstrate that Muse-generated responses enhance people’ capability to correctly identify misinformation by 9.8% (Methods). These high-quality responses and effective corrections by Muse can generally be obtained within minutes of suspicious content appearing on social media (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f). Muse can also inform users of the current lack of evidence, provide an accuracy nudge, and when applicable, express uncertainty (Supp. Fig. S8).
Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (as of February 2024; Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the task’s complexity and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[18] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation. We also changed Muse’s foundation LLM from GPT-4 to open-sourced Llama-3 (70B) and observed two Muses are comparable without modifying any other implementation details (Supp. Fig. S9; Methods), suggesting Muse’s generalizability. Note that the knowledge cutoff for Llama-3 is December 2023, whereas that of GPT-4 is September 2021. Llama-3’s training data might have included some or even all posts used in evaluation and related events, knowledge, and corrections (e.g., their Community Notes data). As a comparison, GPT-4’s training data might only include up to 10% (Methods, where we show that Muse’s performance remains consistent after removing these posts). Hence, it is unfair to directly compare between Llama-3-based and GPT-4-based Muse, and we did not include Llama-3 and its based Muse in evaluation. Nevertheless, our qualitative analysis illustrated in Supp. Fig. S9 reveals the potential to significantly reduce Muse’s cost, as GPT-4 was Muse’s most costly component. Furthermore, as one of the most capable open-sourced LLMs, Llama-3 still cannot accept visual inputs (as of August 2024), emphasizing the importance of Muse’s informative image captioning module.
This work faces several limitations. First, although Muse is capable of responding to misinformation with text and images, it cannot accept video inputs. Second, we only focus on English, one of the most spoken languages in the world. Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has been shown to reduce the spread of misinformation[16] and transparent. X is also a popular social media platform, and one where more than half of users consume news regularly[50] and where misinformation has been shown to diffuse faster than the truth[12]. Fourth, experts assessed and compared Muse against one foundation LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (see Methods and Supp. Fig. S24 for additional results of our ablation study). GPT-4 was chosen as a comparison as it is the state-of-the-art ChatGPT (as of April 2025) and one of the best performing LLMs across a wide range of tasks available today[51, 52].
Acknowledgements
We would like to thank Hacks/Hackers for advertising the study and helping with the recruitment. We also thank members of the UW Social Futures Lab and Behavioral Data Science Group for their suggestions and feedback. This work was supported in part by the National Science Foundation’s Convergence Accelerator program under Award No. 49100421C0037. T.A. and A.S. were supported in part by the Office of Naval Research (#N00014-21-1-2154), NSF IIS-1901386, and NSF CAREER IIS-2142794.
Supplementary material
Supplementary material is available for this paper.
Author contributions
X.Z., A.X.Z., and T.A. conceptualized the work and developed the overarching research goals and aims. X.Z. led the model development. A.S. led the assessment of end-user perceptions of corrections, and X.Z. led the remainder of the evaluation. X.Z. drafted the manuscript, including visualization, and all authors revised the paper. A.X.Z. and T.A. supervised the study.
Data availability
Data, including code, used in this study are available at https://github.com/Social-Futures-Lab/MUSE. We comply with X Terms of Service by only releasing the IDs of tweets. The experts’ names are anonymized.
Supplementary materials
List of supplementary materials
Table LABEL:supp:table:mbfc
Figures S1 to S28
Table S1: Definitions and statistics of factuality and bias categories offered by Media Bias/Fact Check (mediabiasfactcheck.com).
Category Definition # of sources
Factuality:
- Very high: The source is consistently factual, relies on credible information, promptly corrects errors, and has never failed any fact checks in news reporting or opinion pieces. 118
- High: The source is mostly factual and uses mostly credible, low-biased, or high-factual sources. It corrects errors quickly and has failed only one news fact check and up to two op-ed fact checks. 2,313
- Mostly factual: The source is generally accurate but may have a few uncorrected fact-check failures. It can fail up to three op-ed fact checks, especially if it is a low-volume site. While it may use biased sources occasionally, it mostly links to factual content. It is usually pro-science but may sometimes use misleading wording or offer alternative viewpoints. It is reasonably transparent and trustworthy most of the time, but caution is advised. 326
- Mixed: The source may rely on improper sourcing or link to other biased or mixed-factual sources. It often has multiple failed fact checks and does not correct false information or lacks transparency, including the absence of a disclosed mission statement or ownership details. Sources rejecting established scientific consensus on issues such as climate change or vaccines will receive this rating or lower. Sources identified as overt propaganda or designated as hate groups by reputable third-party evaluators will receive this rating or lower due to their inherent bias and potential spread of misleading information. 1,437
- Low: The source is often unreliable and should be fact-checked for fake news, conspiracy theories, and propaganda. 677
- Very low: The source is almost always unreliable and should always be fact-checked for intentional misinformation. 252
Total: 5,123
Bias:
- Least biased: The source has minimal bias and uses very few loaded words (i.e., wording that attempts to influence an audience by using an appeal to emotion or stereotypes). It is factual and usually sourced. 1,054
- Left-center: The source has a slight to moderate liberal bias. It often publishes factual information that utilizes loaded words to favor liberal causes. It is generally trustworthy for information but may require further investigation. 850
- Right-center: Similar to the definition of left-center bias but replacing liberal with conservative. 492
- (Extremely) left: The source is moderately to strongly biased toward liberal causes through story selection or political affiliation. It may utilize strong loaded words, publish misleading reports, and omit reporting of information that may damage liberal causes. It may be untrustworthy. 402
- (Extremely) right: Similar to the definition of (extremely) left bias but replacing liberal with conservative. 314
- Pro-science: The source consists of legitimate science or is evidence-based through the use of credible scientific sourcing. Legitimate science follows the scientific method, is unbiased, and does not use emotional words. The source also respects the consensus of experts in the given scientific field and strives to publish peer-reviewed science. It may have a slight political bias but adheres to scientific principles. 189
- Conspiracy- pseudoscience: The source may publish unverifiable information not always supported by evidence. It may be untrustworthy for credible or verifiable information, so fact-checking and further investigation are recommended on a per-article basis when obtaining information from it. 433
- Questionable: The source exhibits one or more of the following: extreme bias, consistent promotion of propaganda or conspiracies, poor or no sourcing of credible information, a complete lack of transparency, or fake news (i.e., the deliberate attempt to publish hoaxes or disinformation for profit or influence). It may be very untrustworthy and should be fact-checked on a per-article basis. 1,390
- Satire: The source exclusively uses humor, irony, exaggeration, or ridicule to expose and criticize people’s stupidity or vices, particularly in the context of contemporary politics and other topical issues. It does not attempt to deceive. 148
Total: 5,272
Figure S1: Distribution of potential misinformation in X Community Notes (as of February 2023) that received its first response (gray) or first high-quality response (orange) within a certain amount of time.
Figure S2: Examples that show that generating multiple queries helps decompose a post, which may have multiple claims that each needs to be verified, whereas generating one query may overlook some of them and hence lead to not comprehensive identifications of (in)accuracies. Bold text: the verification-needed claims that are overlooked when generating one query but captured when generating more than one query.
Figure S3: Examples that show how retrieved web pages with relatively low relevance to potential misinformation can promote LLM (in this case, GPT-4) hallucinations when generating responses.
Figure S4: Examples of informative image captions, which augment image captions with names of visually represented celebrities and embedded text (see Methods for implementation details).
Figure S5: Examples of Muse-generated responses to accurate, partially accurate, and factually correct but misleading content on social media.
Figure S6: An example of a high-helpfulness response from Community Notes displayed on the corresponding tweet and visible to the public.
Figure S7: Examples of social media content that uses various tactics to make it or part of it false or misleading. These tactics include a: Fabricating a news event (#1), a story with an authentic image (#2, also related to conspiracy theory), and an image with AI (#3). b: Lacking context regarding the motivation behind behavior. Here, it is the consideration that the bill contained no funding for Florida, without which could mislead readers. c: Misinterpreting or misrepresenting a plan, where Biden did not promise to cancel all the debt (#1), or symbol, which is not Nazi $S$ runes but a shorthand for the 46th Separate Airmobile Brigade (#2). d: Using loaded language. e: Improperly analogizing or equating masks blocking respiratory droplets and underwear blocking gas molecules (#1), and COVID-19 and cancer (#2). f: Presenting false, partial, or biased data. 365348 includes multiple reports for the same child and runaways who may not be considered missing because of their guardians know their whereabouts (#1). The data ends in 1885 and is from a specific high elevation site in Greenland, not global temperatures (#2). g: Implying false or oversimplified causation. Many COVID-19 deaths occurred among Medicare beneficiaries because Medicare primarily serves the groups who are at greater risk of adverse outcomes from COVID-19 (#1). The boy was shot when—but not because—he was campaigning for Raphael Warnock (#2). Note that one content may apply more than one tactic intentionally or accidentally.
Figure S8: Examples of Muse-generated responses to social media posts that lack evidence to verify.
Figure S9: Comparison between responses generated by GPT-4, Llama-3, and their based Muse. We observed that GPT4-based Muse and Llama3-based Muse frequently generate similar responses to social media content (see the top example). GPT4-based Muse occasionally generates more comprehensive corrections than Llama3-based Muse (see the middle example). Note that the knowledge cutoff for Llama3 is December 2023, whereas that of GPT-4 is September 2021. In other words, Llama-3’s training data include more recent events and knowledge than GPT-4’s (see the bottom example, where the post was posted in February 2023, and Llama-3 more accurately identifies its falsehood than GPT-4).
Figure S10: Images as examples used for informative image captioning with in-context learning.
   1   {internallinenumbers*} Describe an image in an informative way. Your description should be only based on the given {short caption}, {name of each person}, and {raw text}. If the image is from social media, you should start with "A screenshot of". If the image is a quote from someone, you should start with "A quote from" followed by this person’s name if there is any, then by the quoted text. If the image is an article, you should start with "An article". If the image is a photo, you should start with "A photo of". If the image is a map, you should start with "A map of". {raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.    2   {internallinenumbers*} short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}    3   name of each person: {Fran Lebowitz}    4   raw text: {"In real life, I assure you, there is no such thing as algebra."}    5   image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}    6   short caption: {two men in suits}    7   name of each person: {Jim Caviezel, Michael Emerson}    8   raw text: {}    9   image description: {A photo of Jim Caviezel and Michael Emerson in suits}    10   short caption: {robbie leems on twitter}    11   name of each person: {}    12   {internallinenumbers*} raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother! 1 2 3,790}    13   {internallinenumbers*} image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}    14   short caption: {a moose}    15   name of each person: {}    16   {internallinenumbers*} raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession. finance.yahoo.com Could the stock market power through a recession? ’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}    17   {internallinenumbers*} image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession? ’This would be rare.’" with a picture of a moose. The post was posted at 09:57 22/4/2023.}    18   short caption: {a person pouring tea into a cup}    19   name of each person: {}    20   {internallinenumbers*} raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}    21   {internallinenumbers*} image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup. It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}    22   short caption: {two people standing next to each other with the words love is blind}    23   name of each person: {Nick Lachey}    24   {internallinenumbers*} raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}    25   {internallinenumbers*} image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other. It was written by Hayley Miller and published by MSNBC.}    26   short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}    27   name of each person: {Rahul Kanwal}    28   {internallinenumbers*} raw text: {How engaged are the most-followed journalists on Twitter? Percentage of tweets from each journalist that are at-replies BDUTT 64% RealMikeWilbon 38% 34% ErinAndrews 31% stephenasmith 24% andersoncooper rahulkanwal 13% 13% ninagarcia 5% sardesairajdeep 1% BillSimmons maddow 1% 0% 10% 20% 30% 40% 50% 60% chart by mattmaldre.com}    29   {internallinenumbers*} image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies. The chart was made by mattmaldre.com.}    30   short caption: {a graph showing the global defense budget by region}    31   name of each person: {}    32   {internallinenumbers*} raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}    33   image description: {A graph showing the global defense budget by region. It is from Aviation Week.}    34   short caption: {[IMAGE_CAPTION]}    35   name of each person: {[CELEBRITIES]}    36   raw text: {[OCR]}    37   image description:
Figure S11: LLM prompt for informative image captioning.
Figure S12: Distribution of helpfulness scores of laypeople’s responses in X Community Notes. All: All laypeople’s responses in Community Notes. Average helpfulness: Laypeople’s responses in Community Notes identified with average helpfulness and used in our study. High helpfulness: Laypeople’s responses in Community Notes identified with high helpfulness and used in our study. For $\pm y, xx$: mean, $y$: standard deviation. Community Notes data are regularly updated; ours are up until February 12, 2023.
Figure S13: Distribution of creation times of laypeople’s responses in X Community Notes used in our study. Median of the creation time of laypeople’s average-helpfulness responses: 16 hours after the tweet was posted. Median of the creation time of laypeople’s high-helpfulness responses: 13 hours after the tweet was posted.
Figure S14: Annotation instructions (page 1/7, continued on the next page).
Figure S15: Annotation instructions (page 2/7, continued on the next page).
Figure S16: Annotation instructions (page 3/7, continued on the next page).
Figure S17: Annotation instructions (page 4/7, continued on the next page).
Figure S18: Annotation instructions (page 5/7, continued on the next page).
Figure S19: Annotation instructions (page 6/7, continued on the next page).
Figure S20: Annotation instructions (page 7/7).
Figure S21: Annotation task page.
Figure S22: Impact of starting times of responding to tweets on Muse’s performance. The simulated starting time for Muse (avg): Thirty minutes before the corresponding laypeople’s average-helpfulness responses was created (median: 16 hours after the corresponding tweet was posted; Supplementary Fig. S13). The simulated starting time for Muse (high): Thirty minutes before the corresponding laypeople’s high-helpfulness responses was created (median: 13 hours after the corresponding tweet was posted; Supplementary Fig. S13). The simulated starting time for Muse: The post time of the corresponding tweet (i.e., 0 hours after the corresponding tweet was posted).
Figure S23: Impact of post times of tweets on the performance of Muse and baselines.
(a) Overall quality of responses.
(b) Quality of responses in identifying and explaining (in)accuracies.
(c) Quality of responses in generated text.
(d) Quality of responses in references.
[…]
Figure S25: Impact of popularity of users who posted tweets on Muse’s performance.
Figure S26: Study interface to measure end user perceptions of corrections (page 1/2, continued on the next page).
Figure S27: Study interface to measure end user perceptions of corrections (page 2/2).
Figure S28: We conducted a study on Prolific with 988 participants to assess the real-world impact of Muse and how people perceive and react to AI-generated responses (two-sided Student’s t-test; experiments=16). (a) Only Muse-generated corrections significantly enhanced the correct belief that the misinformation is misleading ($p=0.0001$). (b) None of the corrections made by laypeople, GPT-4, and Muse significantly impacted participants’ intention to share the misinformation, and participants’ intentions were generally low (around 2/7). (c) Muse, GPT-4, and high-helpfulness responses from laypeople all received significantly higher trustworthiness ratings than average-helpfulness responses from laypeople. (d-f) While disclosing the source of the correction (human or AI) had no significant effect on trustworthiness or intention to share, not disclosing the source significantly enhanced the correct belief that the misinformation is misleading ($p=0.0001$).